### v1

# Maintaining Adversarial Robustness in Continuous Learning

Abstract
Adversarial robustness is essential for security and reliability of machine learning systems. However, the adversarial robustness gained by sophisticated defense algorithms is easily erased as the neural network evolves to learn new tasks. This vulnerability can be addressed by fostering a novel capability for neural networks, termed continual robust learning, which focuses on both the (classification) performance and adversarial robustness on previous tasks during continuous learning. To achieve continuous robust learning, we propose an approach called Double Gradient Projection that projects the gradients for weight updates orthogonally onto two crucial subspaces – one for stabilizing the smoothed sample gradients and another for stabilizing the final outputs of the neural network. The experimental results on four benchmarks demonstrate that the proposed approach effectively maintains continuous robustness against strong adversarial attacks, outperforming the baselines formed by combining the existing defense strategies and continual learning methods.
1 Introduction
Continual learning and adversarial robustness are distinct and important research directions in the machine learning (ML) community, each of which has witnessed significant advances. The former addresses a critical challenge known as catastrophic forgetting, where a neural network trained on a new task typically exhibits a dramatic drop in its performance on previously learned tasks if the model cannot revisit the previous data [1]. The latter focuses on developing defenses against adversarial attacks that can deceive models into confidently misclassifying objects by adding subtle targeted perturbations to the input images often imperceptible to human observers [2].
However, the evolution of the neural network’s adversarial robustness in context of continuous learning remains unexplored. Meanwhile, in our experiments, we observe that adversarial robustness gained by well-designed defense algorithms suffers from a phenomenon similar to catastrophic forgetting. Just as in previous work when a dramatic performance drop was observed during continuous learning, an intriguing challenge arises: how can we maintain the adversarial robustness on previous tasks? In other words, we expand the objective in continuous learning to concurrently encompass performance and adversarial robustness.
In this paper, we present a solution inspired by Input Gradients Regularization (IGR) [3] and Gradient Projection Memory (GPM) [4]. With the hypothesis that a neural network with smooth gradients of original samples are more robust against adversarial examples, IGR optimizes neural networks by adding to the objective function a regularization term to penalize gradients of the classification loss with respect to the original samples, which ensures that the predictions of neural networks will not change significantly if any original sample changes locally. At the same time, GPM constraints the back-propagation gradients (used for updating neural network parameters) into the orthogonal direction to the crucial gradient subspace of previous tasks, of which base vectors are formed by performing singular value decomposition on layer-wise inputs of the neural network. GPM can effectively stabilize the final outputs of a neural network, guaranteeing consistent prediction results for data from previous tasks, even after adjusting its weights to accommodate new tasks.
Specifically, we hypothesize that, for a neural network that has smoothed the gradients of samples and is therefore robust for the current task, this robustness will hold if we stabilize these smoothed gradients during continuous learning. Compared to GPM, the crucial gradient subspaces in our approach consist of two parts: one exists in the layer-wise inputs, and the other exists in gradients of these inputs with respect to samples. When learning new tasks, the back-propagation gradients are firstly projected perpendicular to the basis vectors that span the subspaces before being used for weight updates. Our contributions are summarized as follows:
  • 1.
    We introduce a novel capability, referred to as continual robust learning, which is valuable for the adaptive evolution of neural networks across multiple tasks. The key distinction between this and traditional continual learning lies in its focus on not only performance but also adversarial robustness.
  • 2.
    We propose the Double Gradient Projection (DGP) approach, which facilitates continual robust learning by orthogonal projection of back-propagation gradients onto two crucial subspaces, stabilizing both sample gradients and final outputs across previous tasks.
  • 3.
    We demonstrate the effectiveness of our approach on four benchmarks, showcasing its superiority over the baselines established by combining existing algorithms solely for adversarial robustness or continual learning.
2 Background
Figure 1: Feeding data $\mathbf{X}_{p}$ into an exemplar neural network after learning task $\mathcal{T}_{p}$ and $\mathcal{T}_{t}\left(p<t\right)$ respectively. $\Delta\mathbf{W}_{p,t}^{l}$ denotes the change of weights in task $\mathcal{T}_{t}$ relative to task $\mathcal{T}_{p}$. If $\Delta\mathbf{W}_{p,t}^{1}$ meets the constraint $\mathbf{X}_{p}\Delta\mathbf{W}_{p,t}^{1}=0$, then $\mathbf{X}_{p,p}^{2}$ is equal to $\mathbf{X}_{p,t}^{2}$. Recursively, the final outputs $\hat{\mathbf{Y}}_{p,t}$ and $\hat{\mathbf{Y}}_{p,p}$ will be identical even the weights of the neural network are updated. More, if $\Delta\mathbf{W}_{p,t}^{l}$ meets another constraint $\frac{\partial\mathbf{X}^{l}_{p,t}}{\partial\mathbf{X}_{p}}\Delta\mathbf{W}_{p,t}^{l}=0$, the sample gradients $\frac{\partial\hat{\mathbf{Y}}_{p,t}}{\partial\mathbf{X}_{p}}$ and $\frac{\partial\hat{\mathbf{Y}}_{p,p}}{\partial\mathbf{X}_{p}}$ will be identical.
In this section, we introduce the preliminary concepts underlying our work.
2.1 Sample Gradient Regularization
The robustness of the neural network trained with IGR has been demonstrated in adversarial examples across multiple attacks, architectures and datasets [3]. IGR optimizes a neural network $f_{\mathbf{w}}$ by minimizing both the classification loss and the rate of change of that loss with respect to samples, formulated as:
$\mathbf{w}^{\ast}=\underset{\mathbf{w}}{\mathrm{argmin}}H\left(\mathbf{y},\hat{\mathbf{y}}\right)+\lambda\left\|\nabla_{\mathbf{x}}H\left(\mathbf{y},\hat{\mathbf{y}}\right)\right\|,$ (1)
where $H\left(\cdot,\cdot\right)$ is the cross-entropy and $\lambda$ is a hyper-parameter controlling the regular strength. The second term on the right side is to make the variation of the KL divergence between the final output $\hat{\mathbf{y}}$ and the label $\mathbf{y}$ become as small as possible if any sample $\mathbf{x}$ changes locally.
2.2 Gradient Projection Memory
Consider a sequence of task $\left\{\mathcal{T}_{1},\mathcal{T}_{2},\dots\right\}$ where task $\mathcal{T}_{t}$ is associated with paired dataset $\left\{\mathbf{X}_{t},\mathbf{Y}_{t}\right\}$ of size $n_{t}$. When feeding data $\mathbf{X}_{p}$ from previous task $\mathcal{T}_{p}\left(p<t\right)$ into the neural network with optimal weight $\mathbf{W}_{t}$ for task $\mathcal{T}_{t}$ (see Fig. 1), the input and output of the $l$-th linear block (consisting of a linear layer and an activation function $\eta$) are denoted as $\mathbf{X}_{p,t}^{l}$ and $\mathbf{X}_{p,t}^{l+1}$ respectively, then
$\mathbf{X}_{p,t}^{l+1}=\mathbf{X}_{p,t}^{l}\mathbf{W}_{t}^{l}\circ\eta=\mathbf{X}_{p,t}^{l}\left(\mathbf{W}_{p}^{l}+\Delta\mathbf{W}_{p,t}^{l}\right)\circ\eta,$ (2)
where $\Delta\mathbf{W}_{p,t}^{l}$ denotes the change of weights in task $\mathcal{T}_{t}$ relative to task $\mathcal{T}_{p}$. Assuming $\mathbf{X}_{p,t}^{l}=\mathbf{X}_{p,p}^{l}$, a sufficient condition to guarantee $\mathbf{X}_{p,t}^{l+1}=\mathbf{X}_{p,p}^{l+1}$ is by imposing a constraint on $\Delta\mathbf{W}_{p,t}^{l}$ as [5].
$\mathbf{X}_{p,t}^{l}\Delta\mathbf{W}_{p,t}^{l}=0.$ (3)
The final output of a fully-connected network with $L$ linear blocks can be expressed as
$\hat{\mathbf{Y}}_{p,t}=\mathbf{X}_{p}\mathbf{W}_{t}^{1}\circ\eta\circ\mathbf{W}_{t}^{2}\circ\cdots\circ\eta\circ\mathbf{W}_{t}^{L},$ (4)
where $\mathbf{X}_{p,t}^{1}=\mathbf{X}_{p}$. If Eq. 3 is satisfied on each layer recursively, the final outputs $\hat{\mathbf{Y}}_{p,t}$ and $\hat{\mathbf{Y}}_{p,p}$ of the neural network with distinct weights for task $\mathcal{T}_{t}$ and task $\mathcal{T}_{p}$ are identical. Consequently, the performance on task $\mathcal{T}_{p}$ would be maintained after learning task $\mathcal{T}_{t}$.
Figure 2: Graphical representation illustrating the imposed constraints in DGP. (a) The $\mathbf{X}^{l}$ or $\frac{d\mathbf{X}^{l}}{d\mathbf{X}}$ is approximated by $\mathbf{U}^{l}_{k}\mathbf{\Lambda}^{l}_{k}\left(\mathbf{V}^{l}_{k}\right)^{\mathrm{T}}$. (b) Multiplication of $\left(\mathbf{V}^{l}_{k}\right)^{\mathrm{T}}$ with $\Delta\mathbf{W}^{l}$ being zero implies that multiplication of $\mathbf{X}^{l}$ (or $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$) with $\Delta\mathbf{W}^{l}$ is approximately zero. Consequently, weight updates $\Delta\mathbf{W}^{l}$ have little impact on $\mathbf{X}^{l+1}$ (or $\frac{\partial\mathbf{X}^{l+1}}{\partial\mathbf{X}}$) of previous tasks.
GPM [4] performs singular value decomposition (SVD) on $\mathbf{X}_{p,p}^{l}\in\mathbb{R}^{n\times m_{l}}$, where $n$ is the number of samples randomly drawn from the task $\mathcal{T}_{p}$ and $m_{l}$ is the number of features in an input of $l$-th layer:
$\mathbf{X}_{p,p}^{l}\Delta\mathbf{W}_{p,t}^{l}=\mathbf{U}^{l}\mathbf{\Lambda}^{l}\left(\mathbf{V}^{l}\right)^{\mathrm{T}}\Delta\mathbf{W}_{p,t}^{l}\approx\mathbf{U}^{l}_{k}\mathbf{\Lambda}^{l}_{k}\left(\mathbf{V}^{l}_{k}\right)^{\mathrm{T}}\Delta\mathbf{W}_{p,t}^{l},$ (5)
where $\left(\mathbf{V}^{l}\right)^{T}\in\mathbb{R}^{m_{l}\times m_{l}}$ is an orthogonal matrix, of which all the row vectors as a basis span the entire $m_{l}$-dimensional space. Eq. 3 holds true when $\left(\mathbf{V}^{l}\right)^{T}\Delta\mathbf{W}_{p,t}^{l}=0$, indicating that each column vector of $\Delta\mathbf{W}_{p,t}^{l}\in\mathbb{R}^{m_{l}\times m_{l+1}}$ is orthogonal to all the row vectors of $\left(\mathbf{V}^{l}\right)^{T}$. However, it is not possible for a $m_{l}$-dimensional vector to be orthogonal to the entire $m_{l}$-dimensional space unless it is the zero vector, implying no weight update. GPM approximates $\mathbf{X}_{p,p}^{l}$ as $\mathbf{U}^{l}_{k}\mathbf{\Lambda}^{l}_{k}\left(\mathbf{V}^{l}_{k}\right)^{\mathrm{T}}$, where $\left(\mathbf{V}^{l}_{k}\right)^{\mathrm{T}}$ preserves the first $k$ column vectors of $\left(\mathbf{V}^{l}\right)^{\mathrm{T}}$, corresponding to the $k$ largest singular values in diagonal matrix $\mathbf{\Lambda}^{l}$, and spans a subspace of $k\left(<m_{l}\right)$ dimensions. Among all subspaces of $k$ dimensions, weight update orthogonal to this crucial subspace allows for the maximal satisfaction of Eq. 3. An intuitive description is provided in Fig. 2. The value of $k$ is decided by the following criteria:
$\left\|\mathbf{U}^{l}_{k}\mathbf{\Lambda}^{l}_{k}\left(\mathbf{V}^{l}_{k}\right)^{\mathrm{T}}\right\|_{F}^{2}\geq\alpha^{l}\left\|\mathbf{U}^{l}\mathbf{\Lambda}^{l}\left(\mathbf{V}^{l}\right)^{\mathrm{T}}\right\|_{F}^{2},$ (6)
where $\alpha^{l}$ is a given threshold representing the trade-off between learning plasticity and memory stability of the model [6]. By establishing a dedicated pool $\mathcal{P}^{l}$ to retain base vectors $\left(\mathbf{V}^{l}_{k}\right)^{\mathrm{T}}$ from previous tasks, GPM enforces the orthogonality of gradients with respect to these base vectors in the learning process of a new task:
$\nabla_{W^{l}}\mathcal{L}=\nabla_{W^{l}}\mathcal{L}-\left(\nabla_{W^{l}}\mathcal{L}\right)\mathcal{P}^{l}\left(\mathcal{P}^{l}\right)^{T}.$ (7)
For the convolutional layer, the convolution operator can also be formulated as matrix multiplication. Please refer to [7, 4] for details.
3 Method
In this section, we propose an algorithm drawing inspiration from the IGR and GPM to effectively address the challenge of maintaining adversarial robustness in the continuous learning scenario, where revisiting previous data is not feasible. We hypothesize that adversarial robustness, which has been enhanced by IGR in previous tasks, will hold even after learning a sequence of new tasks, if we can stabilize the smoothed gradients of samples from previous tasks.
3.1 Constraint on Weight Updates
By applying the chain rule for derivatives of composite functions, the gradient of the model’s (with $L$ blocks) final output $\hat{\mathbf{y}}$ with respect to a sample $\mathbf{x}$ can be expressed in terms of recursive multiplication:
$\frac{\partial\mathbf{x}^{\left(2\right)}}{\partial\mathbf{x}}\frac{\partial\mathbf{x}^{\left(3\right)}}{\partial\mathbf{x}^{\left(2\right)}}\cdots\frac{\partial\hat{\mathbf{y}}}{\partial\mathbf{x}^{L}}=\frac{\partial\hat{\mathbf{y}}}{\partial\mathbf{x}}.$ (8)
We reformulate Eq. 8 in the Jacobian matrix form
$\begin{bmatrix}\frac{\partial x_{1}^{\left(2\right)}}{\partial x_{1}}&\cdots&\frac{\partial x_{m_{2}}^{\left(2\right)}}{\partial x_{1}}\\ \vdots&\vdots&\vdots\\ \frac{\partial x_{1}^{\left(2\right)}}{\partial x_{m_{1}}}&\cdots&\frac{\partial x_{m_{2}}^{\left(2\right)}}{\partial x_{m_{1}}}\end{bmatrix}\begin{bmatrix}\frac{\partial x_{1}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}\\ \vdots&\vdots&\vdots\\ \frac{\partial x_{1}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}\end{bmatrix}\cdots\\ \begin{bmatrix}\frac{\partial\hat{y}_{1}}{\partial x_{1}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{L}^{l}}\\ \vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{1}}{\partial x_{m_{L}}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{m_{L}}^{L}}\end{bmatrix}=\begin{bmatrix}\frac{\partial\hat{y}_{1}}{\partial x_{1}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{1}}\\ \vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{1}}{\partial x_{m_{1}}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{m_{1}}}\end{bmatrix},$ (9)
where $m_{l}$ represents the number of features in the input of $l$-th block, and $c$ equals the total number of classes within labels.
3.1.1 Linear block
Stringent guarantee. The gradient of the output $\mathbf{x}^{l+1}$ with respect to the input ${\mathbf{x}^{l}}$ of the $l$-th block is derived as explicitly related to the weights $\mathbf{W}^{l}$:
$\frac{\partial\mathbf{x}^{l+1}}{\partial\mathbf{x}^{l}}=\mathbf{W}^{l}\left(\eta^{{}^{\prime}}|\mathbf{x}^{l+1}\right)=\begin{bmatrix}w_{1,1}^{l}&\cdots&w_{m_{l+1},1}^{l}\\ \vdots&\vdots&\vdots\\ w_{1,m_{l}}^{l}&\cdots&w_{m_{l+1},m_{l}}^{l}\end{bmatrix}\begin{bmatrix}\eta^{{}^{\prime}}|{x_{1}^{l+1}}&\cdots&0\\ \vdots&\vdots&\vdots\\ 0&\cdots&\eta^{{}^{\prime}}|{x_{m_{l+1}}^{l+1}}\end{bmatrix}.$ (16)
Each column of the weight matrix (left) represents a single artificial neuron in the linear layer. Element $\eta^{{}^{\prime}}|x_{i}^{l+1}$ in the diagonal matrix (right) represents the derivative of activation function $\eta$ e.g., Relu, of which $\eta^{{}^{\prime}}=1$ if activation $x_{i}^{l+1}>0$, otherwise it is 00. By combining Eq. 8 and Eq. 16, we can efficiently compute the gradient of each block’s input with respect to the sample based on that of the previous block, i.e., $\frac{\partial\mathbf{x}^{l+1}}{\partial\mathbf{x}}=\frac{\partial\mathbf{x}^{l}}{\partial\mathbf{x}}\frac{\partial\mathbf{x}^{l+1}}{\partial\mathbf{x}^{l}}$, instead of having to compute them from scratch which is time-consuming.
We then impose a constraint on weight updates, akin to Eq. 3 but specifically for stabilizing sample gradients (the core idea of this paper):
$\frac{\partial\mathbf{X}^{l}_{p,t}}{\partial\mathbf{X}_{p}}\Delta\mathbf{W}_{p,t}^{l}=0.$ (17)
If Eq. 17 is satisfied on each layer recursively, the sample gradients $\frac{\partial\hat{\mathbf{Y}}_{p,t}}{\partial\mathbf{X}_{p}}$ and $\frac{\partial\hat{\mathbf{Y}}_{p,p}}{\partial\mathbf{X}_{p}}$ of a neural network with distinct weights for task $\mathcal{T}_{t}$ and task $\mathcal{T}_{p}$ are identical (see Fig. 1). Similarly, the method in GPM can be used for an approximate implementation of Eq.17.
Weak guarantee. However, directly performing SVD on the matrix $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}\in\mathbb{R}^{\left(nm_{1}\right)\times m_{l}}$ is computationally time-consuming due to its large size, which is a concat of multiple $\frac{\partial\mathbf{x}^{l}}{\partial\mathbf{x}}\in\mathbb{R}^{m_{1}\times m_{l}}$. To compress the matrix, we modify $\frac{\partial\mathbf{x}^{\left(2\right)}}{\partial\mathbf{x}}$ through column-wise summation, which is located at the beginning of the matrix chain as depicted in Eq. 8, and substitute it back into Eq. 9 as:
$\begin{bmatrix}\sum\limits_{i=1}^{m_{1}}\frac{\partial x_{1}^{\left(2\right)}}{\partial x_{i}}&\cdots\sum\limits_{i=1}^{m_{1}}\frac{\partial x_{m_{2}}^{\left(2\right)}}{\partial x_{i}}\end{bmatrix}\begin{bmatrix}\frac{\partial x_{1}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}\\ \vdots&\vdots&\vdots\\ \frac{\partial x_{1}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}\end{bmatrix}\cdots\\ \begin{bmatrix}\frac{\partial\hat{y}_{1}}{\partial x_{1}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{L}^{l}}\\ \vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{1}}{\partial x_{m_{L}}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{m_{L}}^{L}}\end{bmatrix}=\begin{bmatrix}\sum\limits_{i=1}^{m_{1}}\frac{\partial\hat{y}_{1}}{\partial x_{i}}&\cdots&\sum\limits_{i=1}^{m_{1}}\frac{\partial\hat{y}_{c}}{\partial x_{i}}\end{bmatrix}.$ (18)
According to Eq. 18, $\frac{\partial\mathbf{x}^{l}}{\partial\mathbf{x}}$ transforms to a vector $\in\mathbb{R}^{m_{l}}$. This modification significantly reduces the computational time required for performing SVD on matrix $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}\in\mathbb{R}^{n\times m_{l}}$, while relaxes the stringent guarantee for stabilizing $\frac{\partial\hat{\mathbf{y}}}{\partial\mathbf{x}}$ to a less restrictive one (see right-hand side of Eq. 9 and Eq. 18). The target of the constraint in Eq. 17 is altered from stabilizing the gradient of each final output with respect to each feature in the sample, to stabilizing the sum of gradients of each final output with respect to all features in the sample. This weak guarantee is sufficient to yield desirable results in our experiments for both fully-connected and convolutional neural networks.
3.1.2 Convolutional block
The convolutional block consists of a convolution layer, a batch normalization layer (BN) and an activated function. The gradient of the output $\mathbf{x}^{l+1}$ with respect to an input $\mathbf{x}^{l}$ is derived as:
$\frac{\partial\mathbf{x}^{l+1}}{\partial\mathbf{x}^{l}}=\widetilde{\mathbf{W}}^{l}\partial\mathrm{BN}^{l}\left(\eta^{\prime}\mid\mathbf{x}^{l+1}\right),$ (19)
where $\partial\mathrm{BN}^{l}$ denotes the gradients in BN. The mean and variance $\sigma^{2}$ per-channel used for normalization are constants during evaluation, if they are calculated by tracking during training [8]. In this case, $\partial\mathrm{BN}^{l}$ is a diagonal matrix with the diagonal element $\frac{\gamma}{\sqrt{\sigma^{2}+\epsilon}}$. Please see Appendix A.4 for the case where the mean and variance are batch statistics.
There are two differences between the convolution layer and the linear layer. First, the $\widetilde{\mathbf{W}}^{l}\in\mathbb{R}^{\left(c_{l}h_{l}\omega_{l}\right)\times\left(c_{l+1}h_{l+1}\omega_{l+1}\right)}$ is distinct from the weight matrix $\mathbf{W}^{l}\in\mathbb{R}^{\left(c_{l}k_{l}k_{l}\right)\times c_{l+1}}$, where each column represents a flattened convolution kernel. Here, $c_{l+1}$ ($k_{l}$) denotes the number (size) of kernels in the $l$-th layer, and $h_{l}$ ($\omega_{l}$) denotes the height (width) of the input $\mathbf{x}^{l}$. We give a simple example to illustrate the composition of $\widetilde{\mathbf{W}}^{l}$ in Appendix Fig. 1. The $\widetilde{\mathbf{W}}^{l}$ is sparse, with non-zero elements only present at specific positions of each column, corresponding to the input features that interact with a convolution kernel. To facilitate implementation, we input $\frac{\partial\mathbf{x}^{l}}{\partial\mathbf{x}}$ into the $l$-th convolution layer after reshaping it from $\left(c_{l}h_{l}\omega_{l},\right)$ into $\left(c_{l},h_{l},\omega_{l}\right)$, and subsequently reshape the output from $\left(c_{l+1},h_{l+1},\omega_{l+1}\right)$ back to $\left(c_{l+1}h_{l+1}\omega_{l+1},\right)$, thereby obviating the intricate construction of $\widetilde{\mathbf{W}}^{l}$.
Second, the base vectors formed by directly performing SVD on $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}\in\mathbb{R}^{n\times\left(c_{l}h_{l}\omega_{l}\right)}$ can be used to constrain the updates $\Delta\widetilde{\mathbf{W}}^{l}$ rather than $\Delta\mathbf{W}^{l}$. Consequently, we perform SVD after reshaping $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ into a matrix $\in\mathbb{R}^{\left(nh_{l+1}\omega_{l+1}\right)\times\left(c_{l}k_{l}k_{l}\right)}$ (see Appendix Fig. 2 for details).
3.2 Double Gradient Projection (DGP)
The fundamental principle of our algorithm is concise: stabilizing the smoothed sample gradients (some implementation details are elaborated in the preceding subsection). The overall algorithmic flow is outlined as follows: Firstly, the neural network is trained on task $\mathcal{T}_{t}$ using the loss function that includes a penalty term of sample gradients, as shown in Eq. 1. The weight update is projected to be orthogonal to the base vectors in pool $\mathcal{P}$ if the sequential number $t>1$. Subsequently, after training, SVD is performed on the layer-wise inputs $\mathbf{X}^{l}_{t}$ to obtain base vectors that span the first gradient subspaces crucial for stabilizing the final outputs of the neural network on task $\mathcal{T}_{t}$. Lastly, another SVD is performed on the gradients of the layer-wise inputs with respect to the samples $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}$ to obtain base vectors that span the second gradient subspaces crucial for stabilizing the gradients of final outputs with respect to the samples on task $\mathcal{T}_{t}$. Note that in order to eliminate the redundancy between new bases and existing bases in the pool $\mathcal{P}$, both $\mathbf{X}^{l}_{t}$ and $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}$ are projected orthogonally onto $\mathcal{P}^{l}$ prior to performing SVD. Please refer to [4] for details. A compact pseudo-code of our algorithm is presented in Alg. 1.
Algorithm 1 Double Gradient Projection
Input: Training dataset $\left\{\mathbf{X}_{t},\mathbf{Y}_{t}\right\}$ for task $\mathcal{T}_{t}\in\left\{\mathcal{T}_{1},\mathcal{T}_{2},\dots\right\}$, regularization strength $\lambda$, learning rate $\alpha$
Output: Neural network $f_{w}$ with optimal weights
Initialization: Pool $\mathcal{P}\leftarrow\left\{\right\}$
1:  for  task $\mathcal{T}_{t}\in\left\{\mathcal{T}_{1},\mathcal{T}_{2},\dots\right\}$  do
2:     while  not converged  do
3:        Sample a batch from $\left\{\mathbf{X}_{t},\mathbf{Y}_{t}\right\}$
4:        Calculate $\left\{\nabla_{\mathbf{W}^{l}}\mathcal{L}\right\}$ with the loss $\mathcal{L}$ in Eq. 1
5:        if $t>1$ then
6:           $\nabla_{W^{l}}\mathcal{L}=\nabla_{W^{l}}\mathcal{L}-\left(\nabla_{W^{l}}\mathcal{L}\right)\mathcal{P}^{l}\left(\mathcal{P}^{l}\right)^{T}$ $\forall l=1,2,...,L$triangleright Gradient projection for each layer
7:        end if
8:        $\mathbf{W}^{l}\leftarrow\mathbf{W}^{l}-\alpha\nabla_{\mathbf{W}^{l}}\mathcal{L}$triangleright Weight update
9:     end while
10:     $\mathbf{X}^{l}_{t}\leftarrow\mathbf{X}^{l}_{t}-\mathcal{P}^{l}\left(\mathcal{P}^{l}\right)^{T}\mathbf{X}^{l}_{t}$triangleright Ensure uniqueness of new bases
11:     Perform SVD on $\mathbf{X}^{l}_{t}$ and put bases into $\mathcal{P}^{l}$triangleright Construct first gradient subspace
12:     $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}\leftarrow\frac{\partial\mathbf{X}^{l}_{t}}{\partial\mathbf{X}_{t}}-\mathcal{P}^{l}\left(\mathcal{P}^{l}\right)^{T}\frac{\partial\mathbf{X}^{l}_{t}}{\partial\mathbf{X}_{t}}$
13:     Perform SVD on $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}$ and put bases into $\mathcal{P}^{l}$triangleright Construct second gradient subspace
14:  end for
4 Experiment
4.1 Step
Baselines. For continual learning [9], we adopt (1) SGD, which serves as a naive baseline using stochastic gradient descent to optimize the neural network, (2) EWC [10], (3) SI [11], (4) GEM [12], (5) A-GEM [13], (6) OGD [1] and (7) GPM. For adversarial robustness, we adopt (8) IGR, and (9) Distillation [14].
We combine the methods from fields of continual learning (1$\sim$7) and adversarial robustness (8$\sim$9), such as EWC + IGR, to establish the baselines for continual robust learning. Additionally, we apply the FGSM [15], PGD [16], and AutoAttack [17] to generate adversarial attacks.
Figure 3: ACC varying with the number of learned tasks on datasets of Permuted MNIST (first row), Rotated MNIST (second row), CIFAR100 (third row) and miniImageNet (fourth row). ACC is measured on adversarial samples generated by AutoAttack (first column), PGD (second column) and FGSM (third column), as well as original samples (fourth column). The horizontal axis indicates the number of tasks learned by the neural network at present. The defense strategy used here is IGR. Errors bars denote standard deviation.
Metrics. We use average accuracy (ACC) and backward transfer (BWT) defined as
$\mathrm{ACC}=\frac{1}{T}\sum_{t=1}^{T}R_{T,t},\quad\quad\quad\quad\quad\mathrm{BWT}=\frac{1}{T-1}\sum_{t=1}^{T-1}R_{T,t}-R_{t,t},$ (20)
where $R_{T,t}$ denotes the accuracy of task $t$ at the end of learning task $T$. To evaluate the performance of continuous learning, we measure accuracy on test data from previous tasks. To evaluate the adversarial robustness, we then perturb test data, and re-measure accuracy on the corresponding adversarial samples.
Benchmarks. We evaluate our approach on four supervised benchmarks. Permuted MNIST and Rotated MNIST are variants of MNIST dataset of handwritten digits with 10 tasks applying random permutations of the input pixels and random rotations of the original images respectively [18, 19]. Split-CIFAR100 [11] is a random division of CIFAR100 into 10 tasks, each with 10 different classes. Split miniImageNet is a random division of a subset of the original ImageNet dataset [20] into 10 tasks, each with 5 different classes.
Architectures: The neural network architecture varies across experiments: a fully connected network is used for the MNIST experiments, an AlexNet for the Split-CIFAR100 experiment, and a variant of ResNet18 for the Split-miniImageNet experiment. In both Split-CIFAR100 and Split-miniImageNet experiments, each task has an independent classifier without constraints on weight updates.
The values of $\frac{\partial\mathbf{x}^{2}}{\partial\mathbf{x}}$ (initial term in Eq.16) are solely determined by the weights of the first layer when feeding the same samples (see Fig.1). There are two options to make $\frac{\partial\mathbf{x}^{2}_{p,t}}{\partial\mathbf{x}_{p}}=\frac{\partial\mathbf{x}^{2}_{p,p}}{\partial\mathbf{x}_{p}}$: fixing the first layer after learning task $p$ or assigning an independent first layer to each task. The latter option is chosen for all experiments, as the former seriously diminishes the neural network’s learning ability in subsequent tasks. To ensure fairness, the same setup is applied to the baselines. Further details on architectures can be found in Appendix B.2.
Training details: For the MNIST experiments, the batch size, number of epochs, and input gradient regularization $\lambda$ are set to 32/10/50, respectively. In the Split-CIFAR100 experiments, the values are 10/100/1, and for the Split-miniImageNet experiments, they are 10/20/1. SGD is used as the optimizer. All reported results are averaged over 5 runs with different seeds. The hyperparameter configurations for algorithms of adversarial attack and continuous learning are provided in Appendix B.3.
4.2 Results
4.2.1 Adversarial Robustness
The robustness results on various datasets are presented in left three columns of Fig. 3. Under attacks with different strengths (AutoAttack $>$ PGD $>$ FGSM), the proposed approach (orange lines) consistently exhibits a high level of effectiveness in preserving the robustness of neural networks during continuous learning. In contrast, the baseline such as IGR+GEM (purple lines), which performs well on MNIST datasets against PGD and FGSM attacks, demonstrates a significant decrease when fronted with AutoAttack. The advantage of our approach becomes even more evident when the number of learned tasks increases.
In Fig. 3, the different color lines represent baselines combined by IGR with various methods of continuous learning. Except for the specifically tailored DGP, these methods do not maintain robustness well gained by IGR when the neural network parameters change to learn new tasks. To provide a comprehensive comparison, we include additional experimental results in Fig. 4, focusing on combinations of Distillation with methods of continuous learning. Distillation is another well-known defense method [21] that involves training two models - a teacher model is trained using one-hot ground truth labels and a student model is trained using the softmax probability outputs of the teacher model. The result demonstrates the model’s robustness gained by Distillation is also not effectively maintained against AutoAttack and PGD attacks after learning a sequence of new tasks.
Considering the collective insights presented in Figs. 3 and 4, it is crucial to underscore that the pursuit of an effective defense demands a specially tailored algorithm adept at accommodating variations in model parameters. Mere combinations of existing defense strategies and continual learning methods, as demonstrated in our experiments, fall short of achieving the desired goal of continuous robustness. Our work contributes to the exploration of this under-researched and challenging research direction.
Moreover, there is a notable trend in Fig.4d: the blue line (representing the performance of Distill+GPM on original samples) exhibits a more rapid decline compared to the corresponding blue line in the fourth subplot of the first row in Fig.3 (representing IGR+GPM). Additionally, the purple and blue lines in Fig.4d (representing Distill+GEM and Distill+GPM) closely align with the green line in Fig.4d (representing Distill+SGD). These observations suggest incorporating Distillation into the training procedure compromise the efficacy of these continual learning methods.
Figure 4: As Fig. 3, but for the defense strategy Distillation on PMNIST dataset. Continual learning methods GEM and GPM are selected to combine with Distillation here, which show superior ACC in Fig. 3 compared to other baselines.
4.2.2 Performance
We also assess the ability of the proposed approach for continual learning (ACC on original samples), as illustrated in the fourth column of Fig. 3. Our DGP algorithm demonstrates comparable performance to GPM and GEM on datasets of Permuted MNIST and Rotated MNIST, effectively addressing the issue of catastrophic forgetting on these two datasets. However, results on the datasets of Split-CIFAR indicate that the performance of DGP is slightly inferior to GDP. We speculate that the reason for this could be that DGP stores a larger number of bases after each task than GPM, as DGP constrains the weight updates to be orthogonal to two crucial subspaces – one for stabilizing the final output (required in GPM) and another for stabilizing the sample gradients. Orthogonality to more base vectors restricts weight updates to a narrower subspace, thereby limiting the plasticity of the model. Overall, our approach effectively maintains adversarial robustness while exhibiting continual learning ability.
In addition, it is noteworthy that the performance curves of several classical continual learning methods closely approximate that of naive SGD (green lines). This observation suggest again potential incompatibility between defense strategies (here IGR, i.e., sample gradient penalty) and continuous learning methods. Additional experimental results are provided in Appendix B.1.
4.2.3 Stabilization of sample gradients
Our approach maintains adversarial robustness by stabilizing the smoothed sample gradients. To valid its stabilization effect, we record the variation of sample gradients on the first task during continuous learning process. Specifically, we randomly select $n$ samples at the end of learning $\mathcal{T}_{1}$ and compute their gradients related to correspondingly final outputs. After learning each new task $\mathcal{T}_{t}$, we recompute their gradients. The variation of gradients between $\mathcal{T}_{1}$ and $\mathcal{T}_{t}$ is quantified by similarity measure:
$\mathrm{Sim}=\frac{\mathbf{g}_{1}\mathbf{g}_{t}}{\left|\mathbf{g}_{1}\right|\left|\mathbf{g}_{t}\right|},$ (21)
where $\mathbf{g}$ is a flattened vector of sample gradients. The results on various datasets are presented in Fig. 5. The orange line (representing DGP) shows a relatively flat downward trend, demonstrating the proposed approach indeed has the effect of stabilizing the sample gradients of previous tasks, even after weights of the neural network have been updated.
5 Related Work
Figure 5: Gradient variation of samples from the first task $\mathcal{T}_{1}$ during continuous learning process. The variations are quantified through similarity.
Continual Learning: Optimization. This class of continual learning approaches design and manipulate the optimization programs explicitly, often employing gradient projection as a typical strategy [22]. Replay-based methods like GEM [12] and A-GEM [13] impose constraints on weight updates to align with the direction of experience replay, thereby preserving previous input space and gradient space through historical training samples. Without saving historical training samples, OWM [23] and AOP [24] modify weight updates to the orthogonal direction of the input space of previous tasks. OGD [1] stores the gradient directions of previous tasks and orthogonalizes the current gradient (the gradients of all parameters are combined into a single vector) with respect to them. FS-DGPM [25] dynamically releases unimportant bases of GPM to improve learning plasticity and encourages the convergence towards a flat loss landscape. In contrast to OGD and GPM, Adam-NSCL [5] directly projects candidate weight updates into the unimportant gradient subspace, i.e., the current null space of the uncentered feature covariance from the previous tasks.
Adversarial Robustness: Regularization. As several attack algorithms rely on the model’s gradient to estimate local optima perturbations that can deceive the classifier, a stream of research has focused on adversarial robust optimization by applying regularization approaches to mitigate the impact of small input perturbations on output decisions, such as the spectral norm of the model’s Jacobian matrix [26], the spectral norm of the weight matrix [27] and the generated adversarial perturbation [28].
Our research does not focus on innovation in the single field of continual learning or adversarial robustness, but rather explores a novel direction that requires their integration. It expands the investigation of adversarial robustness to continuous learning scenarios, which holds significant value for enhancing the security and credibility of artificial intelligence systems in their adaptive evolution across multiple tasks or data domains.
6 Limitation and Discussion
In this work, we observe that the adversarial robustness gained by sophisticated defense strategies is easily erased when the neural network learns new knowledge. The direct combinations of existing defense strategies and continuous learning methods fail to effectively address this issue. In the context of continuous learning, a tailored algorithm is necessary to maintain the robustness brought originally by defense strategies. Therefore, we propose Double Gradient Projection, an approach that can prevents the adversarial robustness from rapidly degradation in the face of changes in model parameters.
From the experimental results, our approach has a limitation: as the number of base vectors becomes large, the stability of the neural network is enhanced to hold both robustness and performance across previous tasks. However, the plasticity of the neural network is restricted, which potentially reduces its learning ability for new tasks. Besides, the memory and computational time of our approach are double that of GPM, due to performing SVD twice after each learned task. If there are numerous tasks and the matrix consisting of orthogonal bases reaches full rank, we could approximate this matrix by performing SVD and selecting column vectors corresponding to a fraction of the largest singular values as the new orthogonal bases to free up the rank space.
Appendix A Method
A.1 Sample Gradients
The sample gradients we stabilize in Sec. Method refer to the gradients of the final outputs with respect to samples, rather than gradients of the loss with respect to samples, which are penalized in IGR. Here, we show their relationship:
$\frac{\partial\mathbf{\mathcal{L}}}{\partial\mathbf{x}}=\frac{\partial\mathbf{\mathcal{L}}}{\partial\hat{\mathbf{y}}}\frac{\partial\hat{\mathbf{y}}}{\partial\mathbf{x}}=g\left(\hat{\mathbf{y}}\right)\frac{\partial\hat{\mathbf{y}}}{\partial\mathbf{x}},$ (22)
Where $g$ is a function of $\hat{\mathbf{y}}$. The stabilization of both final outputs $\hat{y}$ and sample gradients $\frac{\partial\hat{\mathbf{y}}}{\partial\mathbf{x}}$ together can result in the stabilization of $\frac{\partial\mathbf{\mathcal{L}}}{\partial\mathbf{x}}$. To maintain the adversarial robustness, achieved by reducing the sensitive of predictions (i.e., final outputs) to subtle changes in samples, it is sufficient to stabilize the smoothed $\frac{\partial\hat{\mathbf{y}}}{\partial\mathbf{x}}$.
A.2 Matrix Composition
A simple example to illustrate the composition of $\widetilde{\mathbf{W}}^{l}$ is depicted in Fig. 6.
Figure 6: Graphic illustration of an example $\widetilde{\mathbf{W}}^{l}$. The shape of an example input $\mathbf{x}^{l}$ and a convolutional kernel $\mathbf{w}^{l}_{i}$ of $l$-th layer is $\left(2,2,2\right)$ and $\left(2,1,1\right)$ respectively. Suppose there are two convolution kernels in total, i.e., $c_{l+1}=2$. The length of each column vector in $\widetilde{\mathbf{W}}^{l}$ is same as the flattened $\mathbf{x}^{l}$, i.e., $c_{l}h_{l}\omega_{l}=2\times 2\times 2=8$. The four subplots in left display the convolution operation of the kernel $\mathbf{w}^{l}_{1}$ on $\mathbf{x}^{l}$, with grey checks indicating the specific input features on which the kernel acts after each slide. The four subplots sequentially correspond to the first four columns of the example $\widetilde{\mathbf{W}}^{l}$. The non-zero elements within each column of $\widetilde{\mathbf{W}}^{l}$ only present (filled by weights of a kernel) at positions corresponding to those specific input features, while the remains are zero-filled.
A.3 Reshape Prior to Performing SVD
A simple example to illustrate why and how to reshape $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ is depicted in Fig. 7.
Figure 7: (a) Performing SVD on an example $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ with the shape $(n,c_{l}h_{l}w_{l})$ obtains the base vectors that can constrain the updates $\Delta\widetilde{\mathbf{W}}^{l}$. Here, $x_{i,j}^{l}$ denotes the $i$-th feature of $j$-th input of $l$-th layer, and an single input $\mathbf{x}^{l}$ from $\mathbf{X}^{l}$ is illustrated on the left of Fig. 6. (b) The orthogonality between any base vector and each $h_{l+1}w_{l+1}(=4)$ column vectors of $\widetilde{\mathbf{W}}^{l}$ (see the right of Fig.6) is equivalent to the orthogonality between $h_{l+1}w_{l+1}$ sub-vectors of the base vector and the weight $\mathbf{w}^{l}$ of the kernel. (c) Prior to performing SVD, each row vector in $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ is reshaped into a matrix consisting of $h_{l+1}w_{l+1}$ row vectors with a length of $c_{l}k_{l}k_{l}$. Consequently, the shape of $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ is modified to $(nh_{l+1}w_{l+1},c_{l}k_{l}k_{l})$.
A.4 Gradients in Batch Normalization Layer
The batch normalization (BN) operation is formalized as
$x_{i}^{out}=\frac{x_{i}^{in}-\mu}{\sqrt{\sigma^{2}+\epsilon}}*\gamma+\beta,$ (23)
Where $\gamma$ and $\beta$ is the learnable weights. When the mean $\mu$ and variance $\sigma^{2}$ per-channel are batch statistics, $x_{i}^{out}$ (feature $i$ in the output of BN) is not only correlated with $x_{i}^{in}$ (feature $i$ in the input of BN), but with the other features of the same channel in the whole batch samples. Therefore, the Jacobian matrix of the $l$-th BN (after $l$-th convolution layer, as shown in Eq. 14 of main text) is a matrix across batch samples with the shape $\left(nc_{l+1}h_{l+1}\omega_{l+1},nc_{l+1}h_{l+1}\omega_{l+1}\right)$, where each element $\frac{\partial x_{i}^{out}}{\partial x_{j}^{in}}$ is given by:
$\begin{cases}\gamma\left[-\left(1-\frac{1}{n}\right)\frac{1}{\sqrt{\sigma^{2}+\epsilon}}-\frac{1}{\left(n-1\right)(\sigma^{2}+\epsilon)^{\frac{3}{2}}}\left(x_{i}-\mu\right)^{2}\right]&\text{ if }i=j,\\ \\ \gamma\left[-\frac{1}{n\sqrt{\sigma^{2}+\epsilon}}-\frac{1}{\left(n-1\right)(\sigma^{2}+\epsilon)^{\frac{3}{2}}}\left(x_{i}-\mu\right)\left(x_{j}-\mu\right)\right]&\text{ if }i\neq j\text{ and }x_{i},x_{j}\text{ are in same channel},\\ \\ 0&\text{ others}.\end{cases}$ (24)
However, due to its extensive scale, the storage or computation of this Jacobian matrix poses high hardware requirements. To facilitate implementation, we propose decomposing this Jacobian matrix into $c_{l+1}$ submatrices with the shape $\left(nh_{l+1}w_{l+1},nh_{l+1}w_{l+1}\right)$, of which all elements belong to the same channel. Subsequently, these submatrices are concatenated to form a new matrix with shape $\left(c_{l+1},nh_{l+1}w_{l+1},nh_{l+1}w_{l+1}\right)$, which effectively optimizes memory usage by eliminating a significant number of zero elements compared to the original Jacobian matrix. Before multiplying with this new matrix, the input gradient matrix of BN should be reshaped from $\left(n,c_{l+1}h_{l+1}w_{l+1}\right)$ to $\left(c_{l+1},nh_{l+1}w_{l+1}\right)$.
Appendix B Experiment
B.1 ACC and BWT on various datasets
The comparisons of ACC and BWT, after learning all the tasks, are presented in Tab. 1 for Permuted MNIST, Tab. 2 for Rotated MNIST and Tab. 3 for Split-CIFAR100 datasets. The majority of baselines exhibit low accuracy (approaching random classification) quickly on the Split-miniImageNet dataset. Therefore, we do not compute BTW for Split-miniImageNet dataset.
Method Permuted MNIST
AutoAttack PGD FGSM Original samples
ACC(%) BWT ACC(%) BWT ACC(%) BWT ACC(%) BWT
SGD 14.1 -0.75 15.4 -0.74 21.8 -0.67 36.8 -0.66
SI 14.3 -0.76 16.5 -0.76 22.3 -0.68 36.9 -0.67
A-GEM 14.1 -0.69 19.7 -0.66 22.9 -0.67 48.4 -0.54
EWC 39.4 -0.47 43.1 -0.48 50.0 -0.35 84.9 -0.12
GEM 12.1 -0.73 75.5 -0.09 72.8 -0.09 96.4 -0.01
OGD 19.7 -0.72 24.1 -0.67 26.0 -0.63 46.8 -0.57
GPM 70.4 -0.11 72.9 -0.10 65.7 -0.12 97.2 -0.01
DGP 81.6 -0.01 81.2 -0.01 75.8 -0.03 97.6 -0.01
Table 1: Comparisons of ACC and BWT after learning all the tasks on the Permuted MNIST dataset.
Method Rotated MNIST
AutoAttack PGD FGSM Original samples
ACC(%) BWT ACC(%) BWT ACC(%) BWT ACC(%) BWT
SGD 14.1 -0.76 9.9 -0.76 20.4 -0.69 32.3 -0.71
SI 13.9 -0.77 15.3 -0.73 20.1 -0.70 33.0 -0.72
A-GEM 14.1 -0.69 21.6 -0.69 24.8 -0.63 45.4 -0.57
EWC 45.1 -0.42 49.5 -0.36 46.5 -0.25 80.7 -0.18
GEM 11.9 -0.73 76.5 -0.08 74.4 -0.08 96.7 -0.01
OGD 19.7 -0.72 23.8 -0.68 23.8 -0.64 48.0 -0.55
GPM 68.8 -0.1 71.5 -0.11 65.9 -0.12 97.1 -0.01
DGP 81.6 0.02 82.6 0.01 78.6 -0.01 98.1 -0.00
Table 2: Comparisons of ACC and BWT after learning all the tasks on the Rotated MNIST dataset.
Method Split-CIFAR100
AutoAttack PGD FGSM Original samples
ACC(%) BWT ACC(%) BWT ACC(%) BWT ACC(%) BWT
SGD 10.3 -0.45 12.8 -0.45 46.5 -0.25 19.4 -0.49
SI 13.0 -0.45 15.2 -0.43 45.4 -0.28 19.8 -0.48
A-GEM 12.6 -0.46 12.9 -0.43 40.6 -0.33 20.7 -0.48
EWC 12.6 -0.43 23.2 -0.31 56.8 -0.15 30.5 -0.35
GEM 21.2 -0.33 19.4 -0.36 60.6 -0.11 47.7 -0.13
OGD 11.8 -0.45 14.1 -0.44 44.2 -0.29 18.9 -0.50
GPM 34.4 -0.13 36.6 -0.17 58.2 -0.16 53.7 -0.10
DGP 36.6 -0.12 39.2 -0.09 67.2 -0.06 48.0 -0.13
Table 3: Comparisons of ACC and BWT after learning all the tasks on the Split-CIFAR100 dataset.
B.2 Architecture Details of Neural Networks
MLP: The fully-connected network in Permuted MNIST and Rotated MNIST experiments consists of three linear layers with 256/256/10 hidden units. No bias units are used. The activation function is Relu. Each task has an independent first layer without constraints imposed on its weight update.
AlexNet: The modified Alexnet[29] in the Split-CIFAR100 experiment consists of three convolutional layers with 32/64/128 kernels of size $\left(4\times 4\right)$/$\left(3\times 3\right)$/$\left(2\times 2\right)$, and three fully connected layers with 2048/2048/10 hidden units. No bias units are used. Each convolution layer is followed by a $\left(2\times 2\right)$ average-pooling layer. The dropout rate is 0.2 for the first two convolutional layers and 0.5 for the remaining layers. The activation function is Relu. Each task has an independent first layer and final layer (classifier) without constraints imposed on its weight update.
ResNet18: The variant ResNet18[30] in the Split-miniImageNet experiment consists of 17 convolutional blocks and one linear layer. The convolutional block comprises a convolutional layer and a batch normalization layer and an Relu activation. The first and last convolutional blocks are followed by a $\left(2\times 2\right)$ average-pooling layer respectively. All convolutional layers use $\left(1\times 1\right)$ zero-padding and kernels of size $\left(3\times 3\right)$. The first convolutional layer has 40 kernels and $\left(2\times 2\right)$ stride, followed by four basic modules, each comprising four convolutional blocks with same number of kernels 40/80/160/320 respectively. The first convolutional layer in each basic modules has $\left(2\times 2\right)$ stride, while the remaining three convolutional layers have $\left(1\times 1\right)$ stride. The skip-connections occur only between basic modules. No bias units are used. In batch normalization layers, tracking mean and variance is used, and the affine parameters are learned in the first task $\mathcal{T}_{1}$, which are then fixed in subsequent tasks. Each task has an independent first layer and final layer.
B.3 Hyper-parameter Configurations
Adversarial attack method method.
dataset Attack method
AutoAttack PGD FGSM
PMNIST $\epsilon=20/255$ $\xi=2/255$,$\delta=40/255$ $\xi=25/255$
RMNIST $\epsilon=20/255$ $\xi=2/255$,$\delta=40/255$ $\xi=25/255$
Split-CIFAR100 $\epsilon=2/255$ $\xi=1/255$,$\delta=4/255$ $\xi=4/255$
Split-miniImageNet $\epsilon=2/255$ $\xi=1/255$,$\delta=4/255$ $\xi=2/255$
Table 4: Hyper-parameter setup for different adversarial attacks
Continuous learning method. We run the methods of SI, EWC, GEM, A-GEM based on the Avalanche [9], an end-to-end continual learning library. In DGP, $\alpha_{1}$ and $\alpha_{2}$ (see $\alpha$ in Eq. 7 of main text) control the number of base vectors added into the pool for stabilizing the final output and sample gradients respectively. $\alpha_{3}$ is used in reducing the number of base vectors when the pool is full (by performing the SVD and $k$-rank approximation on the matrix consisting of all base vectors in the pool).
Dataset Method Hyperparameter
Learning rate Others
Permuted MNIST SGD 0.1 None
SI 0.1 $\lambda=0.1$
EWC 0.1 $\lambda=10$
GEM 0.05 patterns_per_exp $=200$
A-GEM 0.1 sample_size $=64$, patterns_per_exp $=200$
OGD 0.05 memory_size $=300$
GPM 0.05 memory_size $=300$, $\alpha 1=\left[0.95,0.99,0.99\right]$
DGP 0.05
memory_size $=300$, $\alpha 1=[0.95,0.99,0.99]$,
$\alpha 2=0.999$, $\alpha 3=0.996$
Rotated MNIST SGD 0.1 None
SI 0.1 $\lambda=0.1$
EWC 0.1 $\lambda=10$
GEM 0.05 patterns_per_exp $=200$
A-GEM 0.1 sample_size $=64$, patterns_per_exp $=200$
OGD 0.05 memory_size $=300$
GPM 0.05 memory_size $=300$, $\alpha 1=[0.95,0.99,0.99]$
DGP 0.05
memory_size $=300$, $\alpha 1=[0.95,0.99,0.99]$,
$\alpha 2=0.999$, $\alpha 3=0.996$
Split- CIFAR100 SGD 0.05 None
SI $\lambda=0.1$
EWC $\lambda=10$
GEM patterns_per_exp $=200$
A-GEM sample_size $=64$, patterns_per_exp $=200$
OGD memory_size $=300$
GPM memory_size $=100$, $*\alpha 1=0.97+0.003*$task_id
DGP
memory_size $=100$, $*\alpha 1=0.97+0.003*$task_id,
$\alpha 2=0.996$, $\alpha 3=0.99$
Split- miniImageNet SGD 0.1 None
SI $\lambda=0.1$
EWC $\lambda=10$
GEM patterns_per_exp $=200$
A-GEM sample_size $=64$, patterns_per_exp $=200$
OGD memory_size $=100$
GPM memory_size $=100$, $*\alpha 1=0.985+0.003*$task_id
DGP
memory_size $=100$, $\alpha 1=0.96$,
$\alpha 2=0.996$, $\alpha 3=0.996$
Table 5: Hyper-parameter setup in our approach and baselines.

### v2

@@
 Adversarial robustness is essential for security and reliability of machine learning systems.
-However, the adversarial robustness gained by sophisticated defense algorithms is easily erased as the neural network evolves to learn new tasks.
-This vulnerability can be addressed by fostering a novel capability for neural networks, termed continual robust learning, which focuses on both the (classification) performance and adversarial robustness on previous tasks during continuous learning.
-To achieve continuous robust learning, we propose an approach called Double Gradient Projection that projects the gradients for weight updates orthogonally onto two crucial subspaces – one for stabilizing the smoothed sample gradients and another for stabilizing the final outputs of the neural network.
-The experimental results on four benchmarks demonstrate that the proposed approach effectively maintains continuous robustness against strong adversarial attacks, outperforming the baselines formed by combining the existing defense strategies and continual learning methods.
+However, adversarial robustness enhanced by defense algorithms is easily erased as the neural network’s weights update to learn new tasks.
+To address this vulnerability, it is essential to improve the capability of neural networks in terms of robust continual learning.
+Specially, we propose a novel gradient projection technique that effectively stabilizes sample gradients from previous data by orthogonally projecting back-propagation gradients onto a crucial subspace before using them for weight updates.
+This technique can maintaining robustness by collaborating with a class of defense algorithms through sample gradient smoothing.
+The experimental results on four benchmarks including Split-CIFAR100 and Split-miniImageNet, demonstrate that the superiority of the proposed approach in mitigating rapidly degradation of robustness during continual learning even when facing strong adversarial attacks.
 1 Introduction
-Continual learning and adversarial robustness are distinct and important research directions in the machine learning (ML) community, each of which has witnessed significant advances.
-The former addresses a critical challenge known as catastrophic forgetting, where a neural network trained on a new task typically exhibits a dramatic drop in its performance on previously learned tasks if the model cannot revisit the previous data [1].
+Continual learning and adversarial robustness are distinct and important research directions in artificial intelligence, each of which has witnessed significant advances.
+The former addresses a critical challenge known as catastrophic forgetting, where a neural network trained on a sequential of new tasks typically exhibits a dramatic drop in its performance on previously learned tasks if the model cannot revisit the previous data [1].
 The latter focuses on developing defenses against adversarial attacks that can deceive models into confidently misclassifying objects by adding subtle targeted perturbations to the input images often imperceptible to human observers [2].
-However, the evolution of the neural network’s adversarial robustness in context of continuous learning remains unexplored.
-Meanwhile, in our experiments, we observe that adversarial robustness gained by well-designed defense algorithms suffers from a phenomenon similar to catastrophic forgetting.
-Just as in previous work when a dramatic performance drop was observed during continuous learning, an intriguing challenge arises: how can we maintain the adversarial robustness on previous tasks?
-In other words, we expand the objective in continuous learning to concurrently encompass performance and adversarial robustness.
-In this paper, we present a solution inspired by Input Gradients Regularization (IGR) [3] and Gradient Projection Memory (GPM) [4].
-With the hypothesis that a neural network with smooth gradients of original samples are more robust against adversarial examples, IGR optimizes neural networks by adding to the objective function a regularization term to penalize gradients of the classification loss with respect to the original samples, which ensures that the predictions of neural networks will not change significantly if any original sample changes locally.
-At the same time, GPM constraints the back-propagation gradients (used for updating neural network parameters) into the orthogonal direction to the crucial gradient subspace of previous tasks, of which base vectors are formed by performing singular value decomposition on layer-wise inputs of the neural network.
-GPM can effectively stabilize the final outputs of a neural network, guaranteeing consistent prediction results for data from previous tasks, even after adjusting its weights to accommodate new tasks.
-Specifically, we hypothesize that, for a neural network that has smoothed the gradients of samples and is therefore robust for the current task, this robustness will hold if we stabilize these smoothed gradients during continuous learning.
-Compared to GPM, the crucial gradient subspaces in our approach consist of two parts: one exists in the layer-wise inputs, and the other exists in gradients of these inputs with respect to samples.
-When learning new tasks, the back-propagation gradients are firstly projected perpendicular to the basis vectors that span the subspaces before being used for weight updates.
+However, the evolution of the neural network’s adversarial robustness in context of continuous learning remains underexplored.
+In our experiments, we observe that adversarial robustness enhanced by well-designed defense algorithms on previous data is easily lost when the neural network updates its weights to accommodate new tasks, resulting in a phenomenon similar to catastrophic forgetting.
+This presents an intriguing challenge: how can we maintain the adversarial robustness during continuous learning?
+In other words, the objective of continuous learning expands to concurrently encompass (classification) performance and adversarial robustness.
+In this paper, we present a solution by proposing a novel gradient projection technique called Double Gradient Projection (DGP), which inherently enables collaboration with a class of defense algorithms that enhance robustness through sample gradient smoothing.
+DGP is grounded on a theoretical hypothesis that a neural network’s robustness can be maintained if the smoothness of sample gradients from previous data remain unchanged after weight updates.
+Specifically, when learning a new task, DGP projects the back-propagation gradients into the orthogonal direction to a crucial subspace before utilizing them for weight updates.
+This gradient subspace consists of two sets of base vectors derived from previous tasks, which are obtained by performing singular value decomposition on the layer-wise outputs of the neural network and the gradients of layer-wise outputs with respect to samples, respectively.
 Our contributions are summarized as follows:
   • 1.
-    We introduce a novel capability, referred to as continual robust learning, which is valuable for the adaptive evolution of neural networks across multiple tasks.
-The key distinction between this and traditional continual learning lies in its focus on not only performance but also adversarial robustness.
+    We introduce the problem of robust continual learning in the scenario where data from previous tasks cannot be revisited.
   • 2.
-    We propose the Double Gradient Projection (DGP) approach, which facilitates continual robust learning by orthogonal projection of back-propagation gradients onto two crucial subspaces, stabilizing both sample gradients and final outputs across previous tasks.
+    We propose the Double Gradient Projection approach that stabilizes the sample gradients from previous tasks by orthogonally constraining the direction of weight updates.
+It can maintain robustness by collaborating with a class of defense algorithms that enhance robustness through sample gradient smoothing.
   • 3.
-    We demonstrate the effectiveness of our approach on four benchmarks, showcasing its superiority over the baselines established by combining existing algorithms solely for adversarial robustness or continual learning.
-2 Background
+    We validate the superiority of our approach on four image benchmarks.
+Furthermore, the experiment results indicate that without a tailored design, direct combination of existing continual learning and defense algorithms into the training procedure can be conflicting, resulting that the efficacy of the former is seriously weakened.
 Figure 1: Feeding data $\mathbf{X}_{p}$ into an exemplar neural network after learning task $\mathcal{T}_{p}$ and $\mathcal{T}_{t}\left(p<t\right)$ respectively.
@@
 More, if $\Delta\mathbf{W}_{p,t}^{l}$ meets another constraint $\frac{\partial\mathbf{X}^{l}_{p,t}}{\partial\mathbf{X}_{p}}\Delta\mathbf{W}_{p,t}^{l}=0$, the sample gradients $\frac{\partial\hat{\mathbf{Y}}_{p,t}}{\partial\mathbf{X}_{p}}$ and $\frac{\partial\hat{\mathbf{Y}}_{p,p}}{\partial\mathbf{X}_{p}}$ will be identical.
-In this section, we introduce the preliminary concepts underlying our work.
-2.1 Sample Gradient Regularization
-The robustness of the neural network trained with IGR has been demonstrated in adversarial examples across multiple attacks, architectures and datasets [3].
+2 Background
+In this section, we introduce the preliminary concepts underlying our work, including sample gradient smoothing and gradient projection.
+2.1 Sample Gradient Smoothing
+Input gradient regularization (IGR) [3].
+The robustness of the neural network trained with IGR has been demonstrated across multiple attacks, architectures and datasets.
 IGR optimizes a neural network $f_{\mathbf{w}}$ by minimizing both the classification loss and the rate of change of that loss with respect to samples, formulated as:
@@
 The second term on the right side is to make the variation of the KL divergence between the final output $\hat{\mathbf{y}}$ and the label $\mathbf{y}$ become as small as possible if any sample $\mathbf{x}$ changes locally.
-2.2 Gradient Projection Memory
+Adversarial training (AT) [4].
+AT enhances robustness by incorporating adversarial examples generated by Fast Gradient Sign Method (FGSM)  [5] into training data.
+Compared to IGR, which explicitly smooths sample gradients by adding a regularization term into the loss function, AT achieves gradient smoothing implicitly.
+2.2 Gradient Projection
 Consider a sequence of task $\left\{\mathcal{T}_{1},\mathcal{T}_{2},\dots\right\}$ where task $\mathcal{T}_{t}$ is associated with paired dataset $\left\{\mathbf{X}_{t},\mathbf{Y}_{t}\right\}$ of size $n_{t}$.
@@
 where $\Delta\mathbf{W}_{p,t}^{l}$ denotes the change of weights in task $\mathcal{T}_{t}$ relative to task $\mathcal{T}_{p}$.
-Assuming $\mathbf{X}_{p,t}^{l}=\mathbf{X}_{p,p}^{l}$, a sufficient condition to guarantee $\mathbf{X}_{p,t}^{l+1}=\mathbf{X}_{p,p}^{l+1}$ is by imposing a constraint on $\Delta\mathbf{W}_{p,t}^{l}$ as [5].
+Assuming $\mathbf{X}_{p,t}^{l}=\mathbf{X}_{p,p}^{l}$, a sufficient condition to guarantee $\mathbf{X}_{p,t}^{l+1}=\mathbf{X}_{p,p}^{l+1}$ is by imposing a constraint on $\Delta\mathbf{W}_{p,t}^{l}$ as [6, 7]
 $\mathbf{X}_{p,t}^{l}\Delta\mathbf{W}_{p,t}^{l}=0.$ (3)
@@
 Consequently, weight updates $\Delta\mathbf{W}^{l}$ have little impact on $\mathbf{X}^{l+1}$ (or $\frac{\partial\mathbf{X}^{l+1}}{\partial\mathbf{X}}$) of previous tasks.
-GPM [4] performs singular value decomposition (SVD) on $\mathbf{X}_{p,p}^{l}\in\mathbb{R}^{n\times m_{l}}$, where $n$ is the number of samples randomly drawn from the task $\mathcal{T}_{p}$ and $m_{l}$ is the number of features in an input of $l$-th layer:
+Gradient Projection Memory (GPM) [6], designed for improving continual learning ability, performs singular value decomposition (SVD) on $\mathbf{X}_{p,p}^{l}\in\mathbb{R}^{n\times m_{l}}$, where $n$ is the number of samples randomly drawn from the task $\mathcal{T}_{p}$ and $m_{l}$ is the number of features in an input of $l$-th layer:
 $\mathbf{X}_{p,p}^{l}\Delta\mathbf{W}_{p,t}^{l}=\mathbf{U}^{l}\mathbf{\Lambda}^{l}\left(\mathbf{V}^{l}\right)^{\mathrm{T}}\Delta\mathbf{W}_{p,t}^{l}\approx\mathbf{U}^{l}_{k}\mathbf{\Lambda}^{l}_{k}\left(\mathbf{V}^{l}_{k}\right)^{\mathrm{T}}\Delta\mathbf{W}_{p,t}^{l},$ (5)
@@
 $\left\|\mathbf{U}^{l}_{k}\mathbf{\Lambda}^{l}_{k}\left(\mathbf{V}^{l}_{k}\right)^{\mathrm{T}}\right\|_{F}^{2}\geq\alpha^{l}\left\|\mathbf{U}^{l}\mathbf{\Lambda}^{l}\left(\mathbf{V}^{l}\right)^{\mathrm{T}}\right\|_{F}^{2},$ (6)
-where $\alpha^{l}$ is a given threshold representing the trade-off between learning plasticity and memory stability of the model [6].
+where $\alpha^{l}$ is a given threshold representing the trade-off between learning plasticity and memory stability of the neural network [8].
 By establishing a dedicated pool $\mathcal{P}^{l}$ to retain base vectors $\left(\mathbf{V}^{l}_{k}\right)^{\mathrm{T}}$ from previous tasks, GPM enforces the orthogonality of gradients with respect to these base vectors in the learning process of a new task:
@@
 3 Method
-In this section, we propose an algorithm drawing inspiration from the IGR and GPM to effectively address the challenge of maintaining adversarial robustness in the continuous learning scenario, where revisiting previous data is not feasible.
-We hypothesize that adversarial robustness, which has been enhanced by IGR in previous tasks, will hold even after learning a sequence of new tasks, if we can stabilize the smoothed gradients of samples from previous tasks.
+In this section, we propose a novel gradient projection technique inspired from GPM to tackle the challenge of maintaining adversarial robustness in a continuous learning scenario, where revisiting previous data is not feasible.
+We hypothesize that if we can stabilize the sample gradients smoothed by defense algorithms such as IGR and AT on previous tasks, the adversarial robustness of the neural network will hold even after its weights update for learning a sequence of new tasks.
 3.1 Constraint on Weight Updates
-By applying the chain rule for derivatives of composite functions, the gradient of the model’s (with $L$ blocks) final output $\hat{\mathbf{y}}$ with respect to a sample $\mathbf{x}$ can be expressed in terms of recursive multiplication:
+By applying the chain rule for derivatives of composite functions, the gradient of the neural network’s (with $L$ blocks) final output $\hat{\mathbf{y}}$ with respect to a sample $\mathbf{x}$ can be expressed in terms of recursive multiplication:
 $\frac{\partial\mathbf{x}^{\left(2\right)}}{\partial\mathbf{x}}\frac{\partial\mathbf{x}^{\left(3\right)}}{\partial\mathbf{x}^{\left(2\right)}}\cdots\frac{\partial\hat{\mathbf{y}}}{\partial\mathbf{x}^{L}}=\frac{\partial\hat{\mathbf{y}}}{\partial\mathbf{x}}.$ (8)
 We reformulate Eq. 8 in the Jacobian matrix form
-$\begin{bmatrix}\frac{\partial x_{1}^{\left(2\right)}}{\partial x_{1}}&\cdots&\frac{\partial x_{m_{2}}^{\left(2\right)}}{\partial x_{1}}\\ \vdots&\vdots&\vdots\\ \frac{\partial x_{1}^{\left(2\right)}}{\partial x_{m_{1}}}&\cdots&\frac{\partial x_{m_{2}}^{\left(2\right)}}{\partial x_{m_{1}}}\end{bmatrix}\begin{bmatrix}\frac{\partial x_{1}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}\\ \vdots&\vdots&\vdots\\ \frac{\partial x_{1}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}\end{bmatrix}\cdots\\ \begin{bmatrix}\frac{\partial\hat{y}_{1}}{\partial x_{1}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{L}^{l}}\\ \vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{1}}{\partial x_{m_{L}}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{m_{L}}^{L}}\end{bmatrix}=\begin{bmatrix}\frac{\partial\hat{y}_{1}}{\partial x_{1}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{1}}\\ \vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{1}}{\partial x_{m_{1}}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{m_{1}}}\end{bmatrix},$ (9)
+$\begin{bmatrix}\frac{\partial x_{1}^{\left(2\right)}}{\partial x_{1}}&\cdots&\frac{\partial x_{m_{2}}^{\left(2\right)}}{\partial x_{1}}\\ \vdots&\vdots&\vdots\\ \frac{\partial x_{1}^{\left(2\right)}}{\partial x_{m_{1}}}&\cdots&\frac{\partial x_{m_{2}}^{\left(2\right)}}{\partial x_{m_{1}}}\end{bmatrix}\begin{bmatrix}\frac{\partial x_{1}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}\\ \vdots&\vdots&\vdots\\ \frac{\partial x_{1}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}\end{bmatrix}\cdots\\ \begin{bmatrix}\frac{\partial\hat{y}_{1}}{\partial x_{1}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{1}^{L}}\\ \vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{1}}{\partial x_{m_{L}}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{m_{L}}^{L}}\end{bmatrix}=\begin{bmatrix}\frac{\partial\hat{y}_{1}}{\partial x_{1}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{1}}\\ \vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{1}}{\partial x_{m_{1}}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{m_{1}}}\end{bmatrix},$ (9)
 where $m_{l}$ represents the number of features in the input of $l$-th block, and $c$ equals the total number of classes within labels.
@@
 8 and Eq.
-16, we can efficiently compute the gradient of each block’s input with respect to the sample based on that of the previous block, i.e., $\frac{\partial\mathbf{x}^{l+1}}{\partial\mathbf{x}}=\frac{\partial\mathbf{x}^{l}}{\partial\mathbf{x}}\frac{\partial\mathbf{x}^{l+1}}{\partial\mathbf{x}^{l}}$, instead of having to compute them from scratch which is time-consuming.
-We then impose a constraint on weight updates, akin to Eq. 3 but specifically for stabilizing sample gradients (the core idea of this paper):
+10, we can efficiently compute the gradient of each block’s input with respect to the sample based on that of the previous block, i.e., $\frac{\partial\mathbf{x}^{l+1}}{\partial\mathbf{x}}=\frac{\partial\mathbf{x}^{l}}{\partial\mathbf{x}}\frac{\partial\mathbf{x}^{l+1}}{\partial\mathbf{x}^{l}}$, instead of having to compute them from scratch which is time-consuming.
+We then impose a constraint on weight updates for stabilizing sample gradients (the core idea of this work):
 $\frac{\partial\mathbf{X}^{l}_{p,t}}{\partial\mathbf{X}_{p}}\Delta\mathbf{W}_{p,t}^{l}=0.$ (11)
@@
 9 as:
-$\begin{bmatrix}\sum\limits_{i=1}^{m_{1}}\frac{\partial x_{1}^{\left(2\right)}}{\partial x_{i}}&\cdots\sum\limits_{i=1}^{m_{1}}\frac{\partial x_{m_{2}}^{\left(2\right)}}{\partial x_{i}}\end{bmatrix}\begin{bmatrix}\frac{\partial x_{1}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}\\ \vdots&\vdots&\vdots\\ \frac{\partial x_{1}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}\end{bmatrix}\cdots\\ \begin{bmatrix}\frac{\partial\hat{y}_{1}}{\partial x_{1}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{L}^{l}}\\ \vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{1}}{\partial x_{m_{L}}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{m_{L}}^{L}}\end{bmatrix}=\begin{bmatrix}\sum\limits_{i=1}^{m_{1}}\frac{\partial\hat{y}_{1}}{\partial x_{i}}&\cdots&\sum\limits_{i=1}^{m_{1}}\frac{\partial\hat{y}_{c}}{\partial x_{i}}\end{bmatrix}.$ (18)
+$\begin{bmatrix}\sum\limits_{i=1}^{m_{1}}\frac{\partial x_{1}^{\left(2\right)}}{\partial x_{i}}&\cdots\sum\limits_{i=1}^{m_{1}}\frac{\partial x_{m_{2}}^{\left(2\right)}}{\partial x_{i}}\end{bmatrix}\begin{bmatrix}\frac{\partial x_{1}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}\\ \vdots&\vdots&\vdots\\ \frac{\partial x_{1}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}\end{bmatrix}\cdots\\ \begin{bmatrix}\frac{\partial\hat{y}_{1}}{\partial x_{1}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{1}^{L}}\\ \vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{1}}{\partial x_{m_{L}}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{m_{L}}^{L}}\end{bmatrix}=\begin{bmatrix}\sum\limits_{i=1}^{m_{1}}\frac{\partial\hat{y}_{1}}{\partial x_{i}}&\cdots&\sum\limits_{i=1}^{m_{1}}\frac{\partial\hat{y}_{c}}{\partial x_{i}}\end{bmatrix}.$ (12)
 According to Eq.
-18, $\frac{\partial\mathbf{x}^{l}}{\partial\mathbf{x}}$ transforms to a vector $\in\mathbb{R}^{m_{l}}$.
+12, $\frac{\partial\mathbf{x}^{l}}{\partial\mathbf{x}}$ transforms to a vector within the space $\mathbb{R}^{m_{l}}$.
 This modification significantly reduces the computational time required for performing SVD on matrix $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}\in\mathbb{R}^{n\times m_{l}}$, while relaxes the stringent guarantee for stabilizing $\frac{\partial\hat{\mathbf{y}}}{\partial\mathbf{x}}$ to a less restrictive one (see right-hand side of Eq.
 9 and Eq.
-18).
+12).
 The target of the constraint in Eq.
@@
 We give a simple example to illustrate the composition of $\widetilde{\mathbf{W}}^{l}$ in Appendix Fig.
-1.
+6.
 The $\widetilde{\mathbf{W}}^{l}$ is sparse, with non-zero elements only present at specific positions of each column, corresponding to the input features that interact with a convolution kernel.
-To facilitate implementation, we input $\frac{\partial\mathbf{x}^{l}}{\partial\mathbf{x}}$ into the $l$-th convolution layer after reshaping it from $\left(c_{l}h_{l}\omega_{l},\right)$ into $\left(c_{l},h_{l},\omega_{l}\right)$, and subsequently reshape the output from $\left(c_{l+1},h_{l+1},\omega_{l+1}\right)$ back to $\left(c_{l+1}h_{l+1}\omega_{l+1},\right)$, thereby obviating the intricate construction of $\widetilde{\mathbf{W}}^{l}$.
-Second, the base vectors formed by directly performing SVD on $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}\in\mathbb{R}^{n\times\left(c_{l}h_{l}\omega_{l}\right)}$ can be used to constrain the updates $\Delta\widetilde{\mathbf{W}}^{l}$ rather than $\Delta\mathbf{W}^{l}$.
-Consequently, we perform SVD after reshaping $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ into a matrix $\in\mathbb{R}^{\left(nh_{l+1}\omega_{l+1}\right)\times\left(c_{l}k_{l}k_{l}\right)}$ (see Appendix Fig.
+To circumvent the intricate construction of matrix $\widetilde{\mathbf{W}}^{l}$, we identify an alternative approach for implementing $\frac{\partial\mathbf{x}^{l+1}}{\partial\mathbf{x}^{l}}$: reshaping $\frac{\partial\mathbf{x}^{l}}{\partial\mathbf{x}}$ from $\left(c_{l}h_{l}\omega_{l},\right)$ to $\left(c_{l},h_{l},\omega_{l}\right)$, feeding it into the $l$-th convolution layer, and subsequently reshaping the output from $\left(c_{l+1},h_{l+1},\omega_{l+1}\right)$ back to $\left(c_{l+1}h_{l+1}\omega_{l+1},\right)$, i.e., $\frac{\partial\mathbf{x}^{l+1}}{\partial\mathbf{x}}$.
+Second, the base vectors formed by performing SVD on $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}\in\mathbb{R}^{n\times\left(c_{l}h_{l}\omega_{l}\right)}$ (computed through Eq.
+12), can be used directly to constrain the updates $\Delta\widetilde{\mathbf{W}}^{l}$ rather than $\Delta\mathbf{W}^{l}$ (see Appendix Fig.
 7 for details).
+Consequently, we perform SVD after reshaping $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ into a matrix $\in\mathbb{R}^{\left(nh_{l+1}\omega_{l+1}\right)\times\left(c_{l}k_{l}k_{l}\right)}$.
+This results that the base vectors have the same shape $\left(c_{l}k_{l}k_{l},\right)$ with the flattened convolution kernels in the $l$-th layer, and can be used directly to constrain their weight updates.
 3.2 Double Gradient Projection (DGP)
 The fundamental principle of our algorithm is concise: stabilizing the smoothed sample gradients (some implementation details are elaborated in the preceding subsection).
-The overall algorithmic flow is outlined as follows: Firstly, the neural network is trained on task $\mathcal{T}_{t}$ using the loss function that includes a penalty term of sample gradients, as shown in Eq.
-1.
-The weight update is projected to be orthogonal to the base vectors in pool $\mathcal{P}$ if the sequential number $t>1$.
-Subsequently, after training, SVD is performed on the layer-wise inputs $\mathbf{X}^{l}_{t}$ to obtain base vectors that span the first gradient subspaces crucial for stabilizing the final outputs of the neural network on task $\mathcal{T}_{t}$.
-Lastly, another SVD is performed on the gradients of the layer-wise inputs with respect to the samples $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}$ to obtain base vectors that span the second gradient subspaces crucial for stabilizing the gradients of final outputs with respect to the samples on task $\mathcal{T}_{t}$.
+The overall algorithmic flow is outlined as follows: Firstly, the neural network is trained on task $\mathcal{T}_{t}$ with a class of defense algorithms through sample gradient smoothing.
+The weight update is projected to be orthogonal to all the base vectors in pool $\mathcal{P}$ if the sequential number $t>1$.
+Subsequently, after training, SVD is performed on the layer-wise outputs $\mathbf{X}^{l}_{t}$ to obtain base vectors for stabilizing the final outputs of the neural network on task $\mathcal{T}_{t}$.
+Lastly, another SVD is performed on the gradients of the layer-wise outputs with respect to the samples $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}$ to obtain base vectors for stabilizing the gradients of final outputs with respect to the samples on task $\mathcal{T}_{t}$.
 Note that in order to eliminate the redundancy between new bases and existing bases in the pool $\mathcal{P}$, both $\mathbf{X}^{l}_{t}$ and $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}$ are projected orthogonally onto $\mathcal{P}^{l}$ prior to performing SVD.
-Please refer to [4] for details.
 A compact pseudo-code of our algorithm is presented in Alg.
@@
 Algorithm 1 Double Gradient Projection
-Input: Training dataset $\left\{\mathbf{X}_{t},\mathbf{Y}_{t}\right\}$ for task $\mathcal{T}_{t}\in\left\{\mathcal{T}_{1},\mathcal{T}_{2},\dots\right\}$, regularization strength $\lambda$, learning rate $\alpha$
+Input: Training dataset $\left\{\mathbf{X}_{t},\mathbf{Y}_{t}\right\}$ for task $\mathcal{T}_{t}\in\left\{\mathcal{T}_{1},\mathcal{T}_{2},\dots\right\}$, regularization strength $\lambda$ and learning rate $\alpha$
 Output: Neural network $f_{w}$ with optimal weights
@@
 2:     while  not converged  do
-3:        Sample a batch from $\left\{\mathbf{X}_{t},\mathbf{Y}_{t}\right\}$
-4:        Calculate $\left\{\nabla_{\mathbf{W}^{l}}\mathcal{L}\right\}$ with the loss $\mathcal{L}$ in Eq. 1
-5:        if $t>1$ then
-6:           $\nabla_{W^{l}}\mathcal{L}=\nabla_{W^{l}}\mathcal{L}-\left(\nabla_{W^{l}}\mathcal{L}\right)\mathcal{P}^{l}\left(\mathcal{P}^{l}\right)^{T}$ $\forall l=1,2,...,L$triangleright Gradient projection for each layer
-7:        end if
-8:        $\mathbf{W}^{l}\leftarrow\mathbf{W}^{l}-\alpha\nabla_{\mathbf{W}^{l}}\mathcal{L}$triangleright Weight update
-9:     end while
-10:     $\mathbf{X}^{l}_{t}\leftarrow\mathbf{X}^{l}_{t}-\mathcal{P}^{l}\left(\mathcal{P}^{l}\right)^{T}\mathbf{X}^{l}_{t}$triangleright Ensure uniqueness of new bases
-11:     Perform SVD on $\mathbf{X}^{l}_{t}$ and put bases into $\mathcal{P}^{l}$triangleright Construct first gradient subspace
-12:     $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}\leftarrow\frac{\partial\mathbf{X}^{l}_{t}}{\partial\mathbf{X}_{t}}-\mathcal{P}^{l}\left(\mathcal{P}^{l}\right)^{T}\frac{\partial\mathbf{X}^{l}_{t}}{\partial\mathbf{X}_{t}}$
-13:     Perform SVD on $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}$ and put bases into $\mathcal{P}^{l}$triangleright Construct second gradient subspace
-14:  end for
+3:        Sample a batch from $\left\{\mathbf{X}_{t},\mathbf{Y}_{t}\right\}$ and Calculate $\left\{\nabla_{\mathbf{W}^{l}}\mathcal{L}\right\}$
+4:        if $t>1$ then
+5:           $\nabla_{W^{l}}\mathcal{L}=\nabla_{W^{l}}\mathcal{L}-\left(\nabla_{W^{l}}\mathcal{L}\right)\mathcal{P}^{l}\left(\mathcal{P}^{l}\right)^{T}$ $\forall l=1,2,...,L$triangleright Gradient projection for layers
+6:        end if
+7:        $\mathbf{W}^{l}\leftarrow\mathbf{W}^{l}-\alpha\nabla_{\mathbf{W}^{l}}\mathcal{L}$triangleright Weight updates
+8:     end while
+9:     $\mathbf{X}^{l}_{t}\leftarrow\mathbf{X}^{l}_{t}-\mathcal{P}^{l}\left(\mathcal{P}^{l}\right)^{T}\mathbf{X}^{l}_{t}$triangleright Ensure uniqueness for new bases
+10:     Perform SVD on $\mathbf{X}^{l}_{t}$ and put bases into $\mathcal{P}^{l}$triangleright Construct the first set of bases
+11:     $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}\leftarrow\frac{\partial\mathbf{X}^{l}_{t}}{\partial\mathbf{X}_{t}}-\mathcal{P}^{l}\left(\mathcal{P}^{l}\right)^{T}\frac{\partial\mathbf{X}^{l}_{t}}{\partial\mathbf{X}_{t}}$
+12:     Perform SVD on $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}$ and put bases into $\mathcal{P}^{l}$triangleright Construct the second set of bases
+13:  end for
 4 Experiment
@@
 Baselines.
-For continual learning [9], we adopt (1) SGD, which serves as a naive baseline using stochastic gradient descent to optimize the neural network, (2) EWC [10], (3) SI [11], (4) GEM [12], (5) A-GEM [13], (6) OGD [1] and (7) GPM.
-For adversarial robustness, we adopt (8) IGR, and (9) Distillation [14].
-We combine the methods from fields of continual learning (1$\sim$7) and adversarial robustness (8$\sim$9), such as EWC + IGR, to establish the baselines for continual robust learning.
-Additionally, we apply the FGSM [15], PGD [16], and AutoAttack [17] to generate adversarial attacks.
+For continual learning [11], in addition to SGD, which serves as a naive baseline using stochastic gradient descent to optimize the neural network, we adopt six algorithms cover three most important techniques in the field of continual learning: regularization – EWC [12] and SI [13], memory replay – GEM [14] and A-GEM [15], and gradient projection – OGD [1] and GPM [6].
+The fundamental principle of each algorithm are outlined in Appendix B.3.
+For adversarial robustness, we adopt IGR [3] and AT [5].
+We combine algorithms from fields of continual learning and adversarial robustness, such as EWC + IGR, to establish the baselines for robust continual learning.
+On the other hand, we apply the FGSM [5], PGD [16], and AutoAttack [17] to generate adversarial samples.
 Figure 3: ACC varying with the number of learned tasks on datasets of Permuted MNIST (first row), Rotated MNIST (second row), CIFAR100 (third row) and miniImageNet (fourth row).
@@
 The horizontal axis indicates the number of tasks learned by the neural network at present.
-The defense strategy used here is IGR.
+The defense algorithm used here is IGR.
 Errors bars denote standard deviation.
@@
 We evaluate our approach on four supervised benchmarks.
-Permuted MNIST and Rotated MNIST are variants of MNIST dataset of handwritten digits with 10 tasks applying random permutations of the input pixels and random rotations of the original images respectively [18, 19].
-Split-CIFAR100 [11] is a random division of CIFAR100 into 10 tasks, each with 10 different classes.
-Split miniImageNet is a random division of a subset of the original ImageNet dataset [20] into 10 tasks, each with 5 different classes.
+Permuted MNIST and Rotated MNIST are variants of MNIST dataset with 10 tasks applying random permutations of the input pixels and random rotations of the original images respectively [18, 19].
+Split-CIFAR100 [13] is a random division of CIFAR100 into 10 subsets, each with 10 different classes.
+Split-miniImageNet is a random division of a part of the original ImageNet dataset [20] into 10 subsets, each with 5 different classes.
+All images for a specific class are exclusively present in one subset and no overlap of classes between subsets, thus these subsets can be considered as independent datasets, representing a sequence of 10 classification tasks.
 Architectures: The neural network architecture varies across experiments: a fully connected network is used for the MNIST experiments, an AlexNet for the Split-CIFAR100 experiment, and a variant of ResNet18 for the Split-miniImageNet experiment.
@@
 There are two options to make $\frac{\partial\mathbf{x}^{2}_{p,t}}{\partial\mathbf{x}_{p}}=\frac{\partial\mathbf{x}^{2}_{p,p}}{\partial\mathbf{x}_{p}}$: fixing the first layer after learning task $p$ or assigning an independent first layer to each task.
-The latter option is chosen for all experiments, as the former seriously diminishes the neural network’s learning ability in subsequent tasks.
+The latter option is chosen for our experiments, as the former seriously diminishes the neural network’s learning ability in subsequent tasks.
 To ensure fairness, the same setup is applied to the baselines.
@@
 Training details: For the MNIST experiments, the batch size, number of epochs, and input gradient regularization $\lambda$ are set to 32/10/50, respectively.
-In the Split-CIFAR100 experiments, the values are 10/100/1, and for the Split-miniImageNet experiments, they are 10/20/1.
+For the Split-CIFAR100 experiments, the values are 10/100/1, and for the Split-miniImageNet experiments, they are 10/20/1.
 SGD is used as the optimizer.
+The hyperparameter configurations for adversarial attack and continuous learning algorithms are provided in Appendix B.3.
 All reported results are averaged over 5 runs with different seeds.
-The hyperparameter configurations for algorithms of adversarial attack and continuous learning are provided in Appendix B.3.
+We run the experiments on a local machine with three A800 GPUs.
+Figure 4: As Fig.
+3, but for defense algorithm Adversarial Training (AT) on PMNIST dataset.
+Here, we combine AT with continual learning methods GEM and GPM, which have shown superior ACC compared to other baselines in Fig.
+3.
 4.2 Results
 4.2.1 Adversarial Robustness
-The robustness results on various datasets are presented in left three columns of Fig.
-3.
-Under attacks with different strengths (AutoAttack $>$ PGD $>$ FGSM), the proposed approach (orange lines) consistently exhibits a high level of effectiveness in preserving the robustness of neural networks during continuous learning.
+The results about robustness analysis on various datasets are presented in left three columns of Fig.
+3, where different color lines represent the combinations by IGR with diverse continuous learning algorithms and DGP.
+Under attacks with increasing strengths (AutoAttack $>$ PGD $>$ FGSM), the proposed approach (orange lines) consistently exhibits a high level of effectiveness in maintaining robustness of neural networks enhanced by IGR.
 In contrast, the baseline such as IGR+GEM (purple lines), which performs well on MNIST datasets against PGD and FGSM attacks, demonstrates a significant decrease when fronted with AutoAttack.
 The advantage of our approach becomes even more evident when the number of learned tasks increases.
-In Fig.
-3, the different color lines represent baselines combined by IGR with various methods of continuous learning.
-Except for the specifically tailored DGP, these methods do not maintain robustness well gained by IGR when the neural network parameters change to learn new tasks.
-To provide a comprehensive comparison, we include additional experimental results in Fig.
-4, focusing on combinations of Distillation with methods of continuous learning.
-Distillation is another well-known defense method [21] that involves training two models - a teacher model is trained using one-hot ground truth labels and a student model is trained using the softmax probability outputs of the teacher model.
-The result demonstrates the model’s robustness gained by Distillation is also not effectively maintained against AutoAttack and PGD attacks after learning a sequence of new tasks.
+The results of maintaining the robustness enhanced by AT are presented in Fig.
+4.
+The results further demonstrates that baselines fail to effectively maintain the robustness enhanced by AT against AutoAttack and PGD attacks after the neural network learns a sequence of new tasks, whereas GDP performs well.
+Compared to Fig.
+3, the advantage of the proposed method than baselines is more pronounced in Fig.
+4.
 Considering the collective insights presented in Figs.
-3 and 4, it is crucial to underscore that the pursuit of an effective defense demands a specially tailored algorithm adept at accommodating variations in model parameters.
-Mere combinations of existing defense strategies and continual learning methods, as demonstrated in our experiments, fall short of achieving the desired goal of continuous robustness.
-Our work contributes to the exploration of this under-researched and challenging research direction.
-Moreover, there is a notable trend in Fig.4d: the blue line (representing the performance of Distill+GPM on original samples) exhibits a more rapid decline compared to the corresponding blue line in the fourth subplot of the first row in Fig.3 (representing IGR+GPM).
-Additionally, the purple and blue lines in Fig.4d (representing Distill+GEM and Distill+GPM) closely align with the green line in Fig.4d (representing Distill+SGD).
-These observations suggest incorporating Distillation into the training procedure compromise the efficacy of these continual learning methods.
-Figure 4: As Fig.
-3, but for the defense strategy Distillation on PMNIST dataset.
-Continual learning methods GEM and GPM are selected to combine with Distillation here, which show superior ACC in Fig.
-3 compared to other baselines.
-4.2.2 Performance
+3 and 4, it is crucial to underscore that the pursuit of an effective defense demands a tailored algorithm adept at accommodating variations in neural network’s parameters.
+Direct combinations of existing defense strategies and continual learning methods, as demonstrated in our experiments, fall short of achieving the desired goal of continuous robustness.
+4.2.2 Classification Performance
 We also assess the ability of the proposed approach for continual learning (ACC on original samples), as illustrated in the fourth column of Fig.
@@
 Our DGP algorithm demonstrates comparable performance to GPM and GEM on datasets of Permuted MNIST and Rotated MNIST, effectively addressing the issue of catastrophic forgetting on these two datasets.
-However, results on the datasets of Split-CIFAR indicate that the performance of DGP is slightly inferior to GDP.
-We speculate that the reason for this could be that DGP stores a larger number of bases after each task than GPM, as DGP constrains the weight updates to be orthogonal to two crucial subspaces – one for stabilizing the final output (required in GPM) and another for stabilizing the sample gradients.
-Orthogonality to more base vectors restricts weight updates to a narrower subspace, thereby limiting the plasticity of the model.
+However, results on the datasets of Split-CIFAR indicate that the performance of DGP is slightly inferior to GPM.
+We speculate that the reason for this could be that DGP stores a larger number of bases after each task than GPM, as DGP constrains the weight updates to be orthogonal to two sets of base vectors – one for stabilizing the final output (required in GPM) and another for stabilizing the sample gradients.
+Orthogonality to more base vectors restricts weight updates to a narrower subspace, thereby limiting the plasticity of the neural network.
 Overall, our approach effectively maintains adversarial robustness while exhibiting continual learning ability.
-In addition, it is noteworthy that the performance curves of several classical continual learning methods closely approximate that of naive SGD (green lines).
-This observation suggest again potential incompatibility between defense strategies (here IGR, i.e., sample gradient penalty) and continuous learning methods.
-Additional experimental results are provided in Appendix B.1.
+In addition, it is noteworthy that the performance curves of many well-known continual learning algorithms (e.g., EWC) closely approximate that of naive SGD (green lines).
+This important observation suggest a potential incompatibility between existing defense (here IGR) and continuous learning algorithms.
+The effectiveness of the latter can be significantly weakened when they are mixed into the training process.
+For instance, both EWC and IGR add a regularization term into the loss function, but their guidance on the direction of weight updates interferes with each other.
+Experimental results shown the incompatibility with another defense algorithm Distillation [21] are provided in Appendix B.4.
+Figure 5: Gradient variation of samples from the first task $\mathcal{T}_{1}$ during continuous learning process trained with IGR. The variations are quantified through similarity.
 4.2.3 Stabilization of sample gradients
-Our approach maintains adversarial robustness by stabilizing the smoothed sample gradients.
+Our approach maintains adversarial robustness by stabilizing the smoothness of sample gradients.
 To valid its stabilization effect, we record the variation of sample gradients on the first task during continuous learning process.
@@
 5.
-The orange line (representing DGP) shows a relatively flat downward trend, demonstrating the proposed approach indeed has the effect of stabilizing the sample gradients of previous tasks, even after weights of the neural network have been updated.
-5 Related Work
-Figure 5: Gradient variation of samples from the first task $\mathcal{T}_{1}$ during continuous learning process. The variations are quantified through similarity.
-Continual Learning: Optimization.
-This class of continual learning approaches design and manipulate the optimization programs explicitly, often employing gradient projection as a typical strategy [22].
-Replay-based methods like GEM [12] and A-GEM [13] impose constraints on weight updates to align with the direction of experience replay, thereby preserving previous input space and gradient space through historical training samples.
-Without saving historical training samples, OWM [23] and AOP [24] modify weight updates to the orthogonal direction of the input space of previous tasks.
-OGD [1] stores the gradient directions of previous tasks and orthogonalizes the current gradient (the gradients of all parameters are combined into a single vector) with respect to them.
-FS-DGPM [25] dynamically releases unimportant bases of GPM to improve learning plasticity and encourages the convergence towards a flat loss landscape.
-In contrast to OGD and GPM, Adam-NSCL [5] directly projects candidate weight updates into the unimportant gradient subspace, i.e., the current null space of the uncentered feature covariance from the previous tasks.
-Adversarial Robustness: Regularization.
-As several attack algorithms rely on the model’s gradient to estimate local optima perturbations that can deceive the classifier, a stream of research has focused on adversarial robust optimization by applying regularization approaches to mitigate the impact of small input perturbations on output decisions, such as the spectral norm of the model’s Jacobian matrix [26], the spectral norm of the weight matrix [27] and the generated adversarial perturbation [28].
-Our research does not focus on innovation in the single field of continual learning or adversarial robustness, but rather explores a novel direction that requires their integration.
-It expands the investigation of adversarial robustness to continuous learning scenarios, which holds significant value for enhancing the security and credibility of artificial intelligence systems in their adaptive evolution across multiple tasks or data domains.
+The orange line (representing DGP) shows a relatively flat downward trend, demonstrating the proposed approach indeed has the effect of stabilizing the sample gradients of previous tasks, as the neural network’s weights update.
+5 Related Works
+One related work [22] also explores the emerging research direction of robust continual learning.
+A fundamental distinction between that work and ours is their approach requires partial data of previous tasks to be accessible, thereby focusing on the selection of a key subset of previous data and optimizing its utilization.
+In contrast, we follow the stricter yet realistic scenario in the field of continual learning that any data of the previous tasks cannot be revisited.
+Besides, the aim of our work is not to achieve stronger robustness on a single dataset, but rather to maintain robustness across multiple datasets encountered sequentially.
+For further insights into the advancements in adversarial robustness and continual learning, readers are encouraged to refer to dedicated surveys [2, 8].
 6 Limitation and Discussion
-In this work, we observe that the adversarial robustness gained by sophisticated defense strategies is easily erased when the neural network learns new knowledge.
-The direct combinations of existing defense strategies and continuous learning methods fail to effectively address this issue.
-In the context of continuous learning, a tailored algorithm is necessary to maintain the robustness brought originally by defense strategies.
-Therefore, we propose Double Gradient Projection, an approach that can prevents the adversarial robustness from rapidly degradation in the face of changes in model parameters.
-From the experimental results, our approach has a limitation: as the number of base vectors becomes large, the stability of the neural network is enhanced to hold both robustness and performance across previous tasks.
-However, the plasticity of the neural network is restricted, which potentially reduces its learning ability for new tasks.
-Besides, the memory and computational time of our approach are double that of GPM, due to performing SVD twice after each learned task.
-If there are numerous tasks and the matrix consisting of orthogonal bases reaches full rank, we could approximate this matrix by performing SVD and selecting column vectors corresponding to a fraction of the largest singular values as the new orthogonal bases to free up the rank space.
+In this work, we observe that the adversarial robustness gained by well-design defense algorithms is easily erased when the neural network learns new tasks.
+Direct combinations of existing defense and continuous learning algorithms fail to effectively address this issue, and may even give rise to conflicts between them.
+Therefore, we propose a novel gradient projection technique that can mitigate rapidly degradation of robustness in the face of drastic changes in model weights by collaborating with a class of defense algorithms through sample gradient smoothing.
+According to our experiment, the proposed approach has certain limitations.
+First, as the number of base vectors becomes large, the stability of the neural network is enhanced to hold both robustness and performance across previous tasks.
+This stability may restrict the plasticity of the neural network, potentially reducing its ability to learn new tasks.
+Second, if there are numerous tasks and the matrix consisting of orthogonal bases reaches full rank, we approximate this matrix by performing SVD and selecting column vectors corresponding to a fraction of the largest singular values as the new orthogonal bases to free up rank space.
+Three, due to the extra challenge posed by our problem, the perturbation size of adversarial attacks under which the proposed method work effectively is slightly smaller than typical values in adversarial robustness literature (Please see Appendix B.3 for more details).
 Appendix A Method
@@
 The non-zero elements within each column of $\widetilde{\mathbf{W}}^{l}$ only present (filled by weights of a kernel) at positions corresponding to those specific input features, while the remains are zero-filled.
-A.3 Reshape Prior to Performing SVD
+A.3 Reshape $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ Prior to Performing SVD
 A simple example to illustrate why and how to reshape $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ is depicted in Fig. 7.
@@
 Consequently, the shape of $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ is modified to $(nh_{l+1}w_{l+1},c_{l}k_{l}k_{l})$.
+This results that the base vectors obtained from performing SVD on $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ have the same shape $\left(c_{l}k_{l}k_{l},\right)$ with the flattened convolution kernel in the $l$-th layer, and can be directly used to constrain the weight updates of the convolution kernels.
 A.4 Gradients in Batch Normalization Layer
@@
 Before multiplying with this new matrix, the input gradient matrix of BN should be reshaped from $\left(n,c_{l+1}h_{l+1}w_{l+1}\right)$ to $\left(c_{l+1},nh_{l+1}w_{l+1}\right)$.
+A.5 Computational Complexity Increment
+Compared to the naive training procedure, i.e., SGD, the increase of training complexity in the proposed method is mainly related to SVD.
+After finishing the training on each new task, we gather the layer-wise outputs and their gradient with respect to samples, and perform SVD on them to obtain the base vectors used for gradient projection.
+Specifically, we call the interface in Pytorch to perform SVD decomposition.
+This interface uses the Jacobi method with a time complexity of approximately ${o}\left(nm_{l}\min\left(n,m_{l}\right)\right)$, where $n$ is the sample number and $m_{l}$ is the features number of the $l$-th layer’s output.
+Assuming that the neural network consists of $L$ layers, with each layer’s outputs having an equal number of features $m$, the proposed method introduces a computational complexity increment of ${o}\left(Lnm_{l}\min\left(n,m_{l}\right)\right)$.
 Appendix B Experiment
@@
 B.3 Hyper-parameter Configurations
-Adversarial attack method method.
+B.3.1 Adversarial attack algorithm
+The norm of attacks used in experiments are $\ell_{\infty}$. The hyper-parameters in various attack algorithm are provided in Table. 4.
 dataset Attack method
@@
 Split-miniImageNet $\epsilon=2/255$ $\xi=1/255$,$\delta=4/255$ $\xi=2/255$
-Table 4: Hyper-parameter setup for different adversarial attacks
-Continuous learning method.
+Table 4: Hyper-parameter setup to control the attack strength
+The perturbation size used in our experiments are smaller than the typical value in adversarial robustness literature.
+This adjustment is made because when confronted with such intensity of adversarial attacks, regardless of approaches considered (including baselines and the proposed method), the neural network’s robustness on the current task decreases significantly after learning only two or three news tasks.
+Thus, we slightly reduced the perturbation size.
+Then, the advantage of the proposed method becomes evident.
+While most baselines still exhibit a significant decrease after learning only two or three new tasks, the proposed method enables maintaining the model’s robustness after learning a sequence of new tasks.
+B.3.2 Continual learning algorithm
+We outline the fundamental principle for each continual learning algorithm in baselines as follows:
+  • $\bullet$
+    EWC is a regularization technique that utilizes the Fisher Information matrix to quantify the contribution of model parameters on preserving knowledge of previous tasks;
+  • $\bullet$
+    SI computes the local impact of model parameters on global loss variations, consolidating crucial synapses by preventing their modification in new tasks;
+  • $\bullet$
+    A-GEM is a memory-based approach, similar to GEM, which leverages data from episodic memory to adjust the gradient direction of current model update;
+  • $\bullet$
+    OGD is another gradient projection approach where each base vector constrains the weight updates of the entire model, while GPM employs a layer-wise gradient projection strategy.
 We run the methods of SI, EWC, GEM, A-GEM based on the Avalanche [11], an end-to-end continual learning library.
@@
 OGD memory_size $=300$
-GPM memory_size $=100$, $*\alpha 1=0.97+0.003*$task_id
+GPM memory_size $=100$, $\alpha 1=0.97+0.003*$task_id
 DGP
-memory_size $=100$, $*\alpha 1=0.97+0.003*$task_id,
+memory_size $=100$, $\alpha 1=0.97+0.003*$task_id,
 $\alpha 2=0.996$, $\alpha 3=0.99$
@@
 OGD memory_size $=100$
-GPM memory_size $=100$, $*\alpha 1=0.985+0.003*$task_id
+GPM memory_size $=100$, $\alpha 1=0.985+0.003*$task_id
 DGP
@@
 $\alpha 2=0.996$, $\alpha 3=0.996$
-Table 5: Hyper-parameter setup in our approach and baselines.
+Table 5: Hyper-parameter setup in our approach and other CL algorithms in baselines.
+B.4 Incompatibility Between Existing Continual Learning and Defense Algorithms
+Distillation is a well-known defense method [25] that involves training two models - a teacher model is trained using one-hot ground truth labels and a student model is trained using the softmax probability outputs of the teacher model.
+The result of combinations of Distillation and existing continual learning algorithms are presented in Fig.
+8.
+There is a notable trend in Fig.8d: the blue line (representing the performance of Distill+GPM on original samples) exhibits a more rapid decline compared to the corresponding blue line in the fourth subplot of the first row in Fig.3 (representing IGR+GPM), as well as the pink line in Fig.4d (representing AT+GPM).
+Additionally, the purple and blue lines in Fig.8d (representing Distill+GEM and Distill+GPM) closely align with the green line in Fig.8d (representing Distill+SGD).
+These observations suggest again incorporating the defense algorithms, such as Distillation, into the training procedure compromise the efficacy of these continual learning methods.
+Figure 8: As Fig.
+3, but for defense algorithms Distillation on PMNIST dataset.
+Here, we combine Distillation with continual learning methods GEM and GPM, which have shown superior ACC compared to other baselines in Fig.
+3.